### v3

# Large language models are potential rote learners

Large language models are potential rote learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Disentangling Memorization and Capability Learning in Large Language Models
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Are Potentially Rote Memorizer
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Are Both Rote Memorizer and Capacity Learner
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Disentangling Rote Memorization and Capability Learning for
Large Language Model Evaluation
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Large Language Models Could Be Rote Learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
1College of Computer Science and Technology, Zhejiang University
2School of Data Science and Engineering, East China Normal University
3State Key Laboratory of Transvascular Implantation Devices,
The Second Affiliated Hospital Zhejiang University School of Medicine
4School of Public Health, Zhejiang University
5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
6Alibaba Cloud Computing
xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
Abstract
Multiple-choice question (MCQ) benchmarks are widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination. In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle genuine capability acquisition from superficial memorization in LLM evaluation. First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning. To disentangle them, we propose TrinEval, a novel evaluation framework that reformulates MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment. Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
Large Language Models Could Be Rote Learners
Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6 1College of Computer Science and Technology, Zhejiang University 2School of Data Science and Engineering, East China Normal University 3State Key Laboratory of Transvascular Implantation Devices, The Second Affiliated Hospital Zhejiang University School of Medicine 4School of Public Health, Zhejiang University 5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence 6Alibaba Cloud Computing xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
footnotetext: ${}^{\text{\Letter}}$orresponding Author: Haochao Ying.
1 Introduction
The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks (Ouyang et al., 2022; OpenAI, 2024; Touvron et al., 2023). As LLMs continue to improve, evaluating their genuine capacities has emerged as a fundamental challenge, necessitating proper methodologies to ensure fairness and robustness (Ganguli et al., 2023; Liu et al., 2023b).
Among the developed methods, multiple-choice question (MCQ) benchmarks have become a standard approach for evaluation. Typically, LLMs are presented with a question and a fixed set of answer choices, requiring them to select the most appropriate option (see Fig. 1 for illustration). This format enables straightforward performance measurement through accuracy metrics and could cover a wide range of subjects. However, despite their widespread adoption, MCQ-based evaluation raises concerns about reliability due to benchmark contamination (Li and Flanigan, 2024; Kim et al., 2024), i.e., test data unintentionally appears in training corpora and models may exploit memorized content rather than demonstrating genuine understanding, inflating their apparent capabilities. For instance, Zhou et al. (2023) discovers that smaller models with deliberate pre-exposure could outperform their larger counterparts, thereby contradicting widely accepted scaling laws.
Figure 1: MCQ-based LLM evaluation. We observe that LLMs tend to underperform on memorized MCQs.
To mitigate the issue, Zhou et al. (2023) advocates the removal of benchmark datasets from pre-training corpora. However, this strategy conflicts with the fundamental objective of large-scale pre-training, which aims to maximize model performance by exposing LLMs to as much data as possible. From a broader perspective, human learning also involves problem-solving through practicing on similar questions, e.g., exam preparation. While rote memorization of specific questions and answers merely lead to short-term success, repeated practicing can also facilitate deeper conceptual understanding. Inspired, rather than viewing benchmark contamination as a flaw to be eradicated, which is a nearly impossible task at scale (Sainz et al., 2023; Bordt et al., 2024), we argue that it is an inherent aspect of learning and should be accounted for in evaluation. Therefore, this study shifts its focus to evaluating LLMs in the presence of contamination, aiming to distinguish genuine capability gains from superficial memorization effects. The explicit disentangling of these two learning effects remains largely unexplored in MCQ-based evaluation, yet we believe it marks a crucial step towards developing more rigorous and unbiased evaluation methodologies.
To investigate the effects of superficial memorization in LLM evaluation, we compare model performance under different memorization conditions. Inspired by membership inference attacks (MIA) (Carlini et al., 2022a, 2021), we define superficial memorization as an LLM’s ability to verbatim reproduce content, e.g., MCQs in our case. Using this criterion, we partition the MMLU benchmark (Hendrycks et al., 2020) [^1] Selected for its popularity and documented data contamination in widely used LLMs (Sainz et al., 2023). into memorized and non-memorized subsets and evaluate three open-source LLMs [^2] Llama2-7B (Touvron et al., 2023), Mistral-7B-v0.2 (Jiang et al., 2023) and Vicuna-v1.5-7B (Zheng et al., 2023b). on both. Surprisingly, results reveal a consistent yet counterintuitive trend: LLMs perform worse on memorized MCQs than on those not (see Fig. 1 for illustration and Fig. 2 for results). This challenges the assumption that memorization improves model performance and suggests the coexistence of two distinct learning phenomena in LLMs: rote memorization, where models recall content verbatim without true understanding, and genuine capability learning, where they internalize underlying knowledge.
The preliminary investigation has several limitations. First, the binary classification of MCQs as either memorized or non-memorized oversimplifies the nuances of memorization, potentially overlooking intermediate cases. Second, we rely on accuracy to measure performance, which is inherently unreliable. Third, our analysis could not reveal the mutual effects between rote memorization and capability learning. To address these challenges, we propose TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization. TrinEval employs a query-based probing (q-probing) mechanism Allen-Zhu and Li (2023) that reformulates MCQs into an alternative trinity format, i.e., entity-attribute-context. This could prevent direct content recall while preserving knowledge assessment.
Through experiments, we demonstrate that TrinEval’s reformulation is knowledge-preserving, i.e., maintaining testing problems’ inherent knowledge requirements without introducing extra cues, and could effectively reduce memorization. Combined with a continuous superficial memorization quantification metric, TrinEval reveals the in-robustness of LLMs’ capability learning, e.g., with MMLU, tested open-sourced LLMs only mastered 19.6% of knowledge points while 20.5% are memorized by rote in the meanwhile, shedding light on the necessity for further optimization.
2 Related Work
2.1 LLM Evaluation on MCQ Benchmarks
The rapid advancement of LLMs has driven their expansion into diverse domains, necessitating robust and fair evaluation methodologies (Zheng et al., 2023b; Hu et al., 2025) and platforms (Contributors, 2023; Chiang et al., 2024). Among these, evaluating on MCQ benchmarks emerges as a widely adopted approach due to the ease of validation and standardized comparison across models (Hendrycks et al., 2020; Wang et al., 2024; Zhong et al., 2023; Huang et al., 2024).
However, MCQ-based evaluations are not without limitations. Biases in LLM responses have been extensively studied (Dai et al., 2024), revealing issues such as social biases (Salewski et al., 2024; Liu et al., 2023a) and order sensitivity (Akter et al., 2023). To mitigate the latter, Pride (Zheng et al., 2023a) estimates the option positional bias after option permutation. To examine mastery of knowledge, Zhao et al. (2023) applies a hypothesis testing method and checks rephrased-context consistency for a given question. Benchmark contamination is arguably the most severe challenge for MCQ-based evaluations, which may result in misleadingly inflated performance (Zhou et al., 2023; Li and Flanigan, 2024). To address this, prior studies have explored data filtering, frequently-updated test sets (White et al., 2025), and data perturbation (Li et al., 2024).
In this paper, instead of attempting to eliminate contamination, we evaluate LLMs under its presence, aiming to distinguish genuine capability gains from superficial memorization effects. This marks a new perspective of LLM evaluation, revealing the extent to which models truly understand concepts rather than merely memorizing data.
2.2 LLM Memorization
Membership inference attacks (MIA) are commonly used to determine whether a specific sample was present in a model’s training data. Initially studied in smaller models, Carlini et al. (2022b) investigates deep learning memorization mechanisms by identifying and removing easily detectable memorized samples. In the context of LLMs, MIA has been employed to assess privacy risks, revealing that both open- and closed-source models can leak sensitive personal data when provided with related prompts (Kim et al., 2024).
Beyond privacy concerns, Carlini et al. (2022a) formally defines LLM memorization as a model’s ability to verbatim generate text sequences following a prefix prompt. Using this definition, several studies (Sainz et al., 2023; Bordt et al., 2024; Carlini et al., 2021) have examined mainstream LLMs, confirming widespread test data leakage across popular benchmarks. To quantify memorization strength, researchers (Shi et al., 2023; Zhang et al., 2024; Oren et al., 2023; Carlini et al., 2019) have further explored methods such as analyzing token probability distributions in generated outputs. However, while these studies extensively analyze LLM memorization, few explicitly investigate how memorization influences an LLM’s problem-solving ability. In contrast, our work focuses on their interplay, presenting a more rigorous approach to fair and reliable LLM evaluation.
3 Methodology
3.1 Pre-investigation of LLM Capability w.r.t. Memorization
Benchmark contamination often leads to inflated performance estimate. This phenomenon is commonly attributed to models memorizing specific questions and answers rather than demonstrating genuine problem-solving abilities. However, the extent to which and how memorization influences LLM performance remains unclear. To disentangle genuine capability acquisition from superficial memorization, we conduct a preliminary investigation into how LLMs perform under different memorization conditions. By examining model accuracy on memorized vs. non-memorized subsets, we aim to reveal the role of memorization in LLM evaluation and establish a foundation for more rigorous assessment methodologies.
Formally, we define an MCQ as $x=\{x_{Q},x_{O},x_{W}\}$, where $x_{Q}$, $x_{O}$, and $x_{W}$ refer to the question, options, and ground-truth answer, respectively. Following the memorization definition from Carlini et al. (2022a), we say an MCQ $x$ is memorized by LLM $G$ if $G$ can extract/generate the content of options $x_{O}$ exactly given question $x_{Q}$. In practice, we incorporate meta-information (e.g., benchmark name) and 5-shot examples to recall memory and use greedy decoding (i.e., temperature fixed to 0) during extraction (Bordt et al., 2024; Sainz et al., 2023) (refer to Appendix A for the complete prompt). Using MMLU (Hendrycks et al., 2020) as the evaluation benchmark, we divide the test set MCQs into memorized and non-memorized subsets, where the memorized subset consists of 909–982 questions (accounting for 6.5%–7.0% of the total 14,006) depending on the tested LLMs Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B. The detailed statistics of questions across subsets are given in Table 1 of Appendix A, and we also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO (Wang et al., 2024).
Figure 2: Model performance on memorized and non-memorized subsets of MMLU, where ‘0s’ and ‘5s’ stand for zero- and five-shot prompting, respectively.
We then compute the accuracy (ACC) of tested LLMs by subsets as a proxy of model performance under different memorization conditions. The results of both zero- and five-shot prompting are reported in Fig. 2, from which we observe a consistent yet somehow counterintuitive trend: LLMs exhibit 47.2% lower accuracy on average on memorized MCQs compared to non-memorized ones, regardless of LLMs and prompting techniques. This finding challenges the commonly held assumption that memorization directly improves model performance. In addition, it also implies the coexistence of two distinct learning paradigms within LLMs, which we term rote memorization and genuine capability learning, respectively.
However, our pre-investigation has its limitations. The binary classification of memorization potentially overlooks more nuanced forms of learning. Additionally, using ACC as the performance metric does not truly capture model capacity. We address these two issues in the following subsections, which then ensure a disentangle analysis of rote memorization and capability learning.
3.2 Quantifying LLM Memorization
For quantifying the memorization of LLMs, prior research (Shi et al., 2023; Zhang et al., 2024) suggests that outlier tokens, which exhibit higher generation probabilities, are more likely to be found in memorized samples. Building on this idea, we develop a metric that utilizes the bottom $K\%$ of token probabilities within the generated sequence as a measure of memorization. Formally, the memorization score $F_{m}(\overline{x},G)$ of LLM $G$ on text sequence $\overline{x}$ is computed as follows:
$F_{m}(\overline{x},G)=\frac{1}{|\mathcal{M}_{K}(\overline{x})|}\sum\limits_{\overline{x}_{i}\in\mathcal{M}_{K}(\overline{x})}\log p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1}),$ (1)
where $p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1})$ denotes the generation probability of token $\overline{x}_{i}$ by $G$ given its prefix subsequence as context, and set $\mathcal{M}_{K}(\overline{x})$ includes the $K\%$ of tokens with the lowest probabilities. The higher $F_{m}$ is, the more likely $\overline{x}$ is memorized by the LLM, i.e., the least memorized content could still been extracted with a high probability.
3.3 Measuring LLM Capability with TrinEval
We next present TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization.
To understand how LLMs store and manipulate knowledge, Allen-Zhu and Li (2023) created a fictional biography dataset that enumerates various attributes (e.g.,, names, jobs, universities) and trained LLMs on this dataset. They employed a linear query-based probing method to uncover correlations between the entity token embeddings and the associated attributes, revealing that where LLMs encode knowledge, e.g., under person names or sequence of the knowledge mention, is crucial for robust mastery of knowledge. This insight leads us to believe that entity tokens, which should ideally store related knowledge, are the target for evaluating an LLM’s genuine capability.
However, applying this method to real-world datasets, such as MMLU, presents challenges. Unlike controlled datasets with explicitly defined attributes, real-world data includes a far broader range of possible knowledge. As a result, we cannot enumerate all potential attributes and directly apply linear probing. To this end, we propose TrinEval, a verbal query probing method that reformulates MCQs around a knowledge-centric trinity: knowledge entity, attribute, and context. TrinEval is a pluggable augmentation on any MCQ benchmarks and could expose the genuine capability of LLMs by verifying whether they have correctly encoded knowledge. We next explain the elements in the trinity and how to reformulate.
Knowledge entity. We suppose that if an LLM has mastered some knowledge, the key information pertinent to the knowledge should be encoded within a few subject tokens, namely knowledge entity, to support efficient retrieval. By isolating these tokens, TrinEval ensures that only the essential information is considered.
Attribute. The attribute acts as a verbal probe to guide the model focusing on the specific feature or property of the knowledge entity being inquired. This mechanism allows TrinEval to isolate and assess the model’s understanding of the critical aspects of the questioning subject.
Context. In a certain portion of questions, the conditions or background context can significantly influence the solution approach. By explicitly including context in the evaluation process, TrinEval helps the model account for relevant situational details that might otherwise be overlooked, ensuring that the model’s answer is based on a comprehensive understanding of the problem.
By extracting the core and necessary question information in this trinity format, the reformulation by TrinEval is knowledge-preserving for the purposes of assessment. In the meanwhile, it completely destructs the original token sequence, effectively reducing the influence of memorization. We will empirically verify these properties through experiments. The reformulation is completed by a two-round reflection-based prompting method, with detailed procedure (Alg. 1) and related prompts available in Appendix B. Given an MCQ $x=(x_{Q},x_{O},x_{W})$, it first queries a capable reformulation LLM to derive the knowledge entity $x_{E}$, attribute $x_{A}$, and Context $x_{C}$ from the original $x$. The LLM is instructed that the triplet should be sufficient for answering the question correctly, without including the answer option itself, ensuring the integrity of the evaluation. The same LLM then assesses whether the triplet contains all necessary information and no redundant information (typically, the rote-memorization), in the meanwhile, yields a rationale $x_{L}$ as reflection (Shinn et al., 2024; Yao et al., 2022). If it does, the triplet is returned as the re-formulated question. Otherwise, the reformulation model refines the extraction, taking as input $x_{E}$, $x_{A}$, $x_{C}$, and $x_{L}$, and re-evaluates the updated triplet.
Finally, prompting with the extracted $x_{E}$, $x_{A}$, and $x_{C}$ as well as options $x_{O}$, we inspect the generation probability of the ground-truth answer $x_{W}$ as the first token to measure capability:
$F_{c}(x,G)=p_{G}(x_{W}|x_{E},x_{A},x_{C},x_{O}).$ (2)
As can be seen, the $F_{c}$ metric retains the necessary knowledge-centric information while discarding unnecessary biases, especially the rote memorization of LLMs, which leads to the quantification of genuine capability of LLMs.
4 Experiments
In this section, we conduct extensive experiments to answer the following questions:
Q1. Is TrinEval knowledge-preserving in order to fulfill knowledge assessment?
Q2. Can TrinEval reduce memorization effects during capability evaluation?
Q3. What does TrinEval reveal about LLMs’ rote memorization and genuine capability?
4.1 Experiment Setup
Models. We utilize API-based commercial LLMs, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024) for question reformulation by TrinEval. Model evaluation is conducted on open-source LLMs due to limited budgets, and we experiment with three popular LLMs including Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b). All the three LLMs are accessed from Huggingface and implemented with transformers library, we thus could obtain the log-probability of output token for fine-grained study. Throughout our tests, we use the default generation parameters and adopt greedy decoding to enhance reproducibility.
Benchmarks. We evaluate LLM on the widely used MMLU (Hendrycks et al., 2020) benchmark. MMLU consists of 57 subjects from areas including STEM, humanities, social sciences, and others, enabling comprehensive evaluation of LLM capacity. As there are duplicated MCQs across different subjects, we eliminate them and obtain 14,006 MCQs as the test set.
Evaluation. With commercial LLMs, we evaluate model performance by extracting answers with regular expressions. For open-source LLMs, we access the output probability of the first generated token (e.g., option IDs A/B/C/D) to obtain a quantitative performance result.
4.2 Q1. Is TrinEval Knowledge-preserving?
We first verify whether the reformulation by TrinEval is knowledge-preserving in order to fulfill knowledge assessment. To achieve this, our primary objective is to validate that the reformulation approach (1) does not lose key information that results in previous correctly-answered questions being answered incorrectly and (2) does not introduce anomalous or unexpected information that results in inflated performance.
Figure 3: Knowledge-preserving validation for reformulation by TrinEval. 4,343 and 4,645 qualified MCQs are obtained with Qwen and GPT after reformulation. We test Qwen and GPT in these qualified subsets. The green and blue circles stand for the correctly answered MCQs in TrinEval and original formats, respectively.
(a) Evocation with Llama2 and dev-fsp.
(b) Evocation with Mistral and dev-fsp.
(c) Evocation with Vicuan and dev-fsp.
(d) Evocation with Llama2 and seq-fsp.
(e) Evocation with Mistral and seq-fsp.
(f) Evocation with Vicuan and seq-fsp.
Figure 4: The results of memorization evocation under various dataset-related information context, with blue and green curves referring to the memorization difference $\Delta F_{m}$ in the TrinEval and original formats, respectively. In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the training set, and with few-shot prompt from the test set ahead of the current testing question. These curves indicate the growing memorization metric $\Delta F_{m}$ with the stronger dataset-related information in general. However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context are consistently lower than those in the original format, e.g., ‘clean’.
Upon completing the complete TrinEval reformulation process, we ultimately obtained 4,343 MCQs and their corresponding knowledge entities, attributes, and contexts that met our criteria using Qwen, as well as 4,645 qualified MCQs and their respective triplets using GPT. We then instruct Qwen and GPT to answer these respective questions in both the original (baseline) and restated triplet form. The prompts and an MCQ example are available in Table 2 in Appendix and the results are shown in Fig. 3.
We can observe that for Qwen, 92.95% of correctly answered MCQs in the TrinEval format maintain their accuracy in the original format, while for GPT, 90.05% of qualified MCQs are answered correctly, with 95% of these maintaining accuracy in the original format. That is, for both Qwen and GPT, we can infer that the correctly answered MCQs from the qualified ones in TrinEval format can be regarded as a subset of the correctly answered MCQs with the original MCQ format. This proves that the proposed TrinEval reformulation method does not incorporate extra information that leads to additional capability of LLMs. On the other hand, the intersection MCQs between the correctly answered in two formats also make up of around 95% of the MCQs correctly answered in the original MCQ format, which proves that the TrinEval incorporates all the necessary information to answer the question. In conclusion, our TrinEval effectively retains the LLMs’ problem-solving capability compared to the original MCQ text.
4.3 Q2. Can TrinEval Reduce Memorization?
In this subsection, we aim to validate whether the proposed TrinEval can eliminate the unnecessary memorization of LLMs, and thus demonstrate enhanced robustness against various perturbations. To answer this question, following Bordt et al. (2024), we deliberately incorporate the dataset-related information into the context and evaluate whether the TrinEval reformulation can suppress the growing memorization level with memorization evocation of different extent and can reveal the genuine capabilities of LLMs.
We incorporate the dataset-related information into the context, i.e., the name of the dataset, and the few-shot prompt of samples within the same dataset for the memorization-evocation perturbation (see Appendix E). Here we use Llama, Mistral, and Vicuna as the tested LLMs since we access the output probabilities to compute the memorization metric $F_{m}$. As there is no specific zero point of $F_{m}$ indicating the absolute-no memorization of MCQs given an LLM, in order to better visualize the difference between the proposed TrinEval and the original MCQ baseline format, we take the $F_{m}$ with vanilla MCQ (i.e., original MCQ format without any dataset-related prompt) as the baseline and visualize the averaged difference between the $F_{m}$ of the tested format and the baseline.
(a) $F_{m}$ v.s. $F_{c}$ with Llama2 based on Qwen-extracted triplets.
(b) $F_{m}$ v.s. $F_{c}$ with Mistral based on Qwen-extracted triplets.
(c) $F_{m}$ v.s. $F_{c}$ with Vicuna based on Qwen-extracted triplets.
(d) $F_{m}$ v.s. $F_{c}$ with Llama2 based on GPT-extracted triplets.
(e) $F_{m}$ v.s. $F_{c}$ with Mistral based on GPT-extracted triplets.
(f) $F_{m}$ v.s. $F_{c}$ with Vicuna based on GPT-extracted triplets.
Figure 5: The distribution of MCQs based on Memorization metric $F_{m}$ v.s. the Capability metric $F_{c}$. According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs equally into five groups and visualize the distribution of MCQs with the heatmap from weak to strong.
Specifically, for the memorization-evocation permutation, we progressively enhance the prompt context for memory evocation, starting from merely providing the dataset name, to offering samples within the same dataset as few-shot prompts (including the training set of the dataset-dev, and the preceding samples adjacent to the test sample-seq), and finally to providing both as context. Here for each tested MCQ, we calculate the difference between the $F_{m}$ given the corresponding memorization evocation prompt and the $F_{m}$ with the vanilla MCQ baseline. Fig. 4 shows the curve based on the average difference of each MCQ.
From this figure, as stated by Bordt et al. (2024), we can see that $F_{m}$ is growing with the stronger dataset-related context. When providing more specific context related with the test dataset, the LLMs tend to exhibit stronger memorization of the MCQs. Specially, for all three open-source LLMs, the $\Delta F_{m}$ curve of TrinEval is below the curve of the original MCQ baseline. More importantly, the $F_{m}$ of TrinEval with the strongest memorization evocation is still below the vanilla MCQ baseline, which proves that TrinEval can effectively eliminate the memorization from LLMs.
4.4 Q3. TrinEval’s Findings on Memorization and Capability
Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings. ‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$. Further results are shown in Appendix F
In this subsection, we aim to explicitly study the relationship between the memorization and the capability of LLMs with the metrics $F_{m}$ and $F_{c}$. As the commercial-API-based LLMs do not provide the output probability of the whole vocabulary, we mainly use the open-source LLMs to compute these two metrics. After obtaining the $F_{m}$ and $F_{c}$ of each MCQ, we separate all the qualified MCQs into 5 equal groups. Finally, we utilize the heatmap to reveal the relationship between the capability and the memorization of the tested LLMs.
As shown in Fig. 5, most of the MCQs concentrate on the lower left corner and the upper right corner of the heatmap. Specifically, for $F_{m}$ v.s. $F_{c}$ with Llama2 based on Qwen-extracted triplets, MCQs within the lower left $2\times 2$ squares and the upper right $2\times 2$ squares make up of the 38.57% of all the tested MCQs with a Pearson-correlation of -0.7755 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 74.17% of all the tested MCQs with a Pearson-correlation of -0.8124 (p-value $<0.05$). For the results with Mistral based on Qwen-extracted triplets, MCQs within the lower left and upper right $2\times 2$ squares make up of the 44.90% of all the tested MCQs with a Pearson-correlation of -0.8722 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 80.82% of all the tested MCQs with a Pearson-correlation of -0.8794 (p-value $<0.05$). More results are shown in Tab. 3. This evidence indicates that MCQs with lower memorization levels tend to exhibit better problem-solving capabilities of LLMs, while those with higher memorization levels are associated with reduced performance in solving tasks.
Next, we hypothesize that the LLMs are potential rote learners through the human memory system, which has been characterized by two fundamental components: Long-Term Memorization (LTM) and Short-Term Memorization (STM) Shiffrin (2003). Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability. In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework. This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts. As illustrated in Allen-Zhu and Li (2023) and Ovadia et al. (2023), LLMs trained with multiple rephrased corpus tend to perform better than LLMs trained with only the original corpus. When providing only one format of training corpus, similar to the STM system, LLMs tend to memorize the corpus at token-level rather than knowledge-level. In other words, LLMs encode these corpora at a shallow level with the original format. After questions are rephrased with methods like our proposed TrinEval, the input corpus seems connected with the known knowledge like the LTM for structured storage and enables sophisticated information generalization. We show more detailed results in Appendix C.
To further validate our hypothesis, we compute the embeddings of MCQs within the qualified MMLU dataset and average the distance between the other closest $1\%$ MCQs. We visualize the mean distance of MCQs within the lower left and upper right $2\times 2$ squares in Fig. 5. The results are shown in Fig. 6. We surprisingly find that the averaged distance of the Genuine Capability Learning MCQs (i.e., MCQs within the upper right $2\times 2$ squares) is almost half as much as the distance of the Rote Memorization MCQs (i.e., MCQs within the lower left $2\times 2$ squares). The result hints that the memorized MCQs are sparsely encoded by MCQs while the non-memorized ones share common embeddings, which is again coincident with the findings of the STM and LTM.
Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.
5 Conclusion
This study provided a novel perspective on benchmark contamination in LLM evaluation, reframing it as an inherent aspect of learning. This perspective led us to explore the relationship between memorization and genuine capability in LLMs. Through our empirical investigation, we observed a surprising result: LLMs performed worse on memorized MCQs compared to those not, suggesting that superficial memorization may undermine problem-solving ability rather than enhance it. This finding also implies the existence of two distinct learning paradigms in LLMs: rote memorization and genuine capability learning.
To disentangle them, we proposed TrinEval, a novel evaluation method that reformulates MCQs into a knowledge-centric trinity, thus separating the influence of memorization from genuine knowledge application. Experiments validated both the knowledge-preserving and memorization-reducing properties of this approach. Based on that, TrinEval reveals the in-robustness of LLMs’ knowledge learning, e.g., popular open-source LLMs memorize 20.5% of knowledge points by rote without understanding in MMLU. As such, we believe this work lays the groundwork for future studies on improving LLM knowledge robustness and more thorough evaluation.
6 Limitations
Our limitations are mainly two points. First, though our proposed TrinEval retrains the problem-solving ability of the LLMs and obtains stronger robustness, it is not a dynamical re-organizing method that can still be leaked and pre-experienced during training. On the one hand, we appeal to the LLM developers not to use this re-organizing method as part of the training corpus. On the other hand, future works will be focused on developing dynamic evaluation method (Zhu et al., 2023, 2024). Second, we did not give a clear exploration on how and why the more LLMs memorize, the less the capability of the LLMs obtains. In future work, we will also look into the mechanism of the training and structure of LLMs for a thorough study of the phenomenon.
Appendix A Details of the extracting prompts and the extracted (memorized) MCQs
In this section, we introduce the details of the processed dataset and the prompts for extraction. MCQs from some subjects contain similar or identical options [^3] E.g., the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’).. With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio. In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset. After deduplication, we obtain 14,006 MCQs for evaluation. The extraction prompt and the detailed statistics are shown in the following text and Tab. 1.
Prompt template for extraction:
You are an expert of multiple choice questions of MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A.
Prompt template for pre-investigation on LLM Memorization w.r.t. Capability:
You are an expert of multiple choice questions of MMLU dataset. The following are multiple choice questions (with answers) about [subject]. [examples] [question] Options: A. [content for option A] B. [content for option B] C. [content for option C] D. [content for option D] Answer:
{tblr}
cells = c, cell21 = r=3, cell51 = r=3, cell81 = r=3, vline2-3 = 1-2,5,8, vline2-3 = 3-4,6-7,9-10, hline1-2,5,8,11 = -, Model & Subset Simple Pro MMLU
Llama memorized 912 70 982
non-mem. 6,548 6,476 13,024
all 7,460 6,546 14,006
Mistral memorized 879 36 915
non-mem. 6,581 6,510 13,091
all 7,460 6,546 14,006
Vicuna memorized 893 16 909
non-mem. 6,567 6,530 13,097
all 7,460 6,546 14,006
Table 1: Statistics of memorized and non-memorized questions by Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B in MMLU.
Appendix B Details of TrinEval
In this section, we introduce the details of the proposed TrinEval. The pseudo-code is shown in the Alg. 1. The prompts used are also shown below. Note that the potential data leakage is often caused by the data crawled on the Huggingface dataset site. Thus, we also provide the original text of MCQs in the format on the Huggingface dataset site to mimic the data contamination with in-context learning.
Algorithm 1 MCQ reformulation by TrinEval
0:  Question $x_{Q}$, options $x_{O}$, and answer $x_{W}$ of an MCQ.
0:  Reformulated question $x^{R}_{Q}$.
1:  Preliminarily extract knowledge entity $x_{E}$, attribute $x_{A}$, and context $x_{C}$ based on $x_{Q}$, $x_{O}$ and $x_{W}$;
2:  Initialize $X^{R}_{Q}={x_{E},x_{A},x_{C}}$;
3:  Validate the adequacy and necessity of the $x^{R}_{Q}$ and give reasons $x_{L}$;
4:  if $x^{R}_{Q}$ matches the requirement then
5:     Return $x^{R}_{Q}$;
6:  else
7:     Re-extract $x_{E}^{\prime}$, $x_{A}^{\prime}$, and $x_{C}^{\prime}$ by reflecting with $x_{E},x_{A},x_{C}$ and $x_{L}$;
8:     Update $x^{R}_{Q}={x_{E}^{\prime},x_{A}^{\prime},x_{C}^{\prime}}$;
9:     Validate the adequacy and necessity of the $x^{R}_{Q}$ and give reasons $x_{L}$;
10:     if $x^{R}_{Q}$ matches the requirement then
11:        Return $x^{R}_{Q}$;
12:     else
13:        Discard the MCQ, return $None$;
14:     end if
15:  end if
Prompt template for triplet extraction:
You are an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to obtain the key information of the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] # Notes - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should be the original text within the Question and should not incorporate any phrases that cannot be exactly matched in the Question. - Never include any information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Prompt template for triplet validation & reflection:
You are an expert of [subject] and an advanced reasoning agent that can determine whether the given Knowledge Keyword, Attribute of the Knowledge Keyword and the Context present most of the necessary information of the Question for obtaining the given Answer. Suppose you have sufficient background knowledge about subj. Consider the given Knowledge Keyword, Attribute and the Context, then determine whether the given Answer can be directly obtained from them even without the Question. # Steps 1. **Check the Semantic completeness:** Suppose you have sufficient background knowledge about [subject], and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Check if the given Knowledge Keyword, Attribute, Context are the original text within the Question and contain the necessary queried information the Question itself provided (ignore the information the Question did not provided). If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 2. **Check the Answer relevance:** Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Read through the given Knowledge Keyword, Attribute, Context and the given Answer. Check if the Answer can be directly inferred with the given Knowledge Keyword, Attribute and the Context without seeing the Question. If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 3. **Check the Semantic Redundancy:** Read through the given Knowledge Keyword, Attribute, Context, the given Question and the given corresponding Answer. Check if the Answer can be directly matched within the given Knowledge Keyword, Attribute and the Context. Check if there are any unnecessary information within the given Knowledge Keyword, Attribute and the Context for obtaining the given Answer to the Question. If not so, point out what is redundant. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Verdict for the given Knowledge Keyword, Attribute and Context:** [Single verdict (Yes/No) here for whether the given Knowledge Keyword, Attribute and Context contain most of the asked information of the Question, can be used to infer the given Answer with only them without the whole Question, and do not contain redundant information for obtaining the given Answer.] # Notes - Do not deviate from the specified format. Do not generate anything else after the Verdict (only Yes/No) for the given Knowledge Keyword, Attribute and Context. - Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. For Semantic completeness and Answer relevance, it is acceptable to miss information that is also not incorporated in the Question. - Provide a detailed explanation following the given steps before arriving at the verdict (Yes/No). Provide a final verdict (only Yes/No) in order at the end in the given format. - **Question:** [question] - **Answer:** [answer] - **Knowledge Keyword:** [extracted knowledge entity] - **Attribute:** [extracted attribute] - **Context:** [extracted context]
Prompt template for the second round triplet extraction:
You are an advanced reasoning agent that can improve through self-reflection and an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to rephrase the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] You will be given a previous trial. You were unsuccessful in extracting the Knowledge Keyword, Attribute and the necessary that meet the requirements in the previous trial. Given the Reflection below, improve the process. The process is as follows: # Previous returns: - **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] - **Question:** [question] - **Answer:** [answer option content] - **Knowledge Keyword:** [extracted knowledge entity of the last trial] - **Attribute:** [attribute of the last trial] - **Context:** [context of the last trial] - **Reflection:** [rational of the last trial] # Notes - Consider the Reflection given above. Improve the extraction of Knowledge Keyword, Attribute and Context (if necessary). - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword should be phrases within the Question and should not incorporate any information of the Fact corpus or the given Answer that is not mentioned in the Question. - The extracted Attribute and Context (if necessary) should only include information from the Question corpus. Never include information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Original MCQ TrinEval MCQ
You are an expert on multiple choice questions of [subject]. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** [question] - **Options:** A. [option A] B. [option B] C. [option C] D. [option D] You are an expert on multiple choice questions of [subject]. Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question. Determine the correct answer option to the question. The Knowledge Entity is the questioned subject of the question. The Attribute is the questioned attribute of the Knowledge Entity, and the Context is the necessary context information for answering the question. Given a set of Knowledge Entity, Attribute, and Context (which three are extracted as the key information from a question), and the potential Answer options to the Question, analyze the given Knowledge Entity, Attribute, Context as well as the options. Generate the option ID of the correct option (answer). - **Knowledge Entity:** [knwoledge entity] - **Attribute:** [attribute] - **Context:** [context] - **Options:** A. [option A] B. [option B] C. [option C] D. [option D]
Original MCQ Example TrinEval MCQ Example
You are an expert on multiple choice questions of high school computer science. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** Which of the following is usually NOT represented in a subroutine’s activation record frame for a stack-based programming language? - **Options:** A. Values of local variables B. A heap area C. The return address D. Stack pointer for the calling activation record You are an expert on multiple choice questions of high school computer science. Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question. Determine the correct answer option to the question. The Knowledge Entity is the questioned subject of the question. The Attribute is the questioned attribute of the Knowledge Entity, and the Context is the necessary context information for answering the question. Given a set of Knowledge Entity, Attribute, and Context (which three are extracted as the key information from a question), and the potential Answer options to the Question, analyze the given Knowledge Entity, Attribute, Context as well as the options. Generate the option ID of the correct option (answer). - **Knowledge Entity:** subroutine’s activation record frame - **Attribute:** usually NOT represented - **Context:** for a stack-based programming language - **Options:** A. Values of local variables B. A heap area C. The return address D. Stack pointer for the calling activation record
Table 2: Template and an example of the Original MCQ template and the TrinEval MCQ template. [$\cdot$] refers to the blank that should be filled according to the content of each MCQ.
Appendix C Detailed results of memorization v.s. capability
In this section, we exhibit the detailed results of the Q3. What does TrinEval reveal about the memorization v.s. the capability of LLMs. We reveal the ratio of MCQs within the upper right and lower left 2 $\times$ 2 and 3 $\times$ 3 squares as well as the Pearson correlations between the $F_{m}$ and $F_{c}$ of these MCQs. Our analysis reveals a tendency towards a negative correlation between the capabilities and memorization of LLMs shown in the Tab. 3.
(a) Probability and p-value with Llama2 based on GPT.
(b) Probability and p-value with Mistral based on GPT.
(c) Probability and p-value with Vicuna based on GPT.
(d) Probability and p-value with Llama2 based on GPT (21.59 % MCQs filtered).
(e) Probability and p-value with Mistral based on GPT (15.95 % MCQs filtered).
(f) Probability and p-value with Vicuna based on GPT (17.59 % MCQs filtered).
Figure 7: The over-performing probability curve and p-value curve with different $F_{m}$ thresholds. In this figure, we take each unique $F_{m}$ as the threshold to separate the qualified MCQs as the Memorized and Capable MCQs. We compute the probability of a randomly selected Capable MCQ’s $F_{c}$ exceeds a randomly selected Memorized MCQ’s $F_{c}$ under each threshold as the blue curve, and the green curve is the p-value of the T-test between the $F_{c}$s of the Capable MCQs and the Memorized MCQs.
Further, inspired by the Precision-Recall Curve, we take each unique $F_{m}$ of the qualified MCQs as the threshold to separate them as the Memorized and Capable MCQs. For each separation, we compute the probability of whether the $F_{c}$ of a randomly selected Capable MCQ exceeds the $F_{c}$ of a randomly selected Memorized MCQ and plot them as the blue curve. We also compute the T-test p-value between the $F_{c}$s of the Memorized MCQs and Capable MCQs as the green curve. The results are shown in Fig. 7. For the second row, we filter out the MCQs within the upper left and lower right $2\times 2$ squares. From the figure, we observe that over a relatively long segment in the middle of the x-axis threshold range, the probability remains at a comparatively high value, while the p-value stays below 0.05. From this, we can conclude that $F_{m}$ can distinguish between MCQs with high $F_{c}$ and those with low $F_{c}$ with a negative correlation at a high confidence level. This further supports that LLMs are potential rote learners, the more the LLMs memorize, the more poorly they perform.
{tblr}
row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=2c, cell15 = c=2c, cell31 = r=3, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell61 = r=3, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell91 = r=3, cell93 = c, cell94 = c, cell95 = c, cell96 = c, cell103 = c, cell104 = c, cell105 = c, cell106 = c, cell113 = c, cell114 = c, cell115 = c, cell116 = c, cell121 = r=3, cell123 = c, cell124 = c, cell125 = c, cell126 = c, cell133 = c, cell134 = c, cell135 = c, cell136 = c, cell143 = c, cell144 = c, cell145 = c, cell146 = c, cell151 = r=3, cell153 = c, cell154 = c, cell155 = c, cell156 = c, cell163 = c, cell164 = c, cell165 = c, cell166 = c, cell173 = c, cell174 = c, cell175 = c, cell176 = c, cell181 = r=3, cell183 = c, cell184 = c, cell185 = c, cell186 = c, cell193 = c, cell194 = c, cell195 = c, cell196 = c, cell203 = c, cell204 = c, cell205 = c, cell206 = c, hline1,3,21 = -, hline2 = 3-6, LLMs & Dataset $2\times 2$ squares $3\times 3$ squres
Ratio (%) Pearson correlation Ratio (%) Pearson correlation
Llama2-Qwen All 38.57 -0.7755 74.17 -0.8124
Simple 37.07 -0.7784 72.63 -0.8121
Pro 38.66 -0.783 74.51 -0.8109
Llama2-GPT All 35.22 -0.7835 71.04 -0.7924
Simple 33.9 -0.777 69.62 -0.7919
Pro 35.45 -0.7916 71.54 -0.7881
Mistral-Qwen All 44.9 -0.8722 80.82 -0.8794
Simple 38.47 -0.8494 74.04 -0.8271
Pro 44.32 -0.8045 80.08 -0.8682
Mistral-GPT All 40.37 -0.8042 76.58 -0.8736
Simple 35.51 -0.8297 72.27 -0.8664
Pro 38.52 -0.7103 74.91 -0.7969
Vicuna-Qwen All 42.94 -0.8771 79.23 -0.8365
Simple 37.86 -0.758 73.85 -0.7168
Pro 42.01 -0.8609 77.86 -0.886
Vicuna-GPT All 38.69 -0.8621 74.83 -0.8672
Simple 34.77 -0.8096 70.71 -0.7775
Pro 37.37 -0.7794 73.98 -0.8728
Table 3: The ratio and the Pearson-correlation between the $F_{c}$ and $F_{m}$ of the MCQs within the upper right and lower left $2\times 2$ and $3\times 3$ squares. For LLMs, ‘Llama2-Qwen’ refers that the $F_{c}$ and $F_{m}$ are calculated with Llama2 based on the Qwen-extracted triplet, and similarly hereinafter. For the Dataset column, ‘All’ stands for all the qualified MCQs after the triplet extraction, ‘Pro’ refers to the qualified MCQs that are the members of the mmlupro dataset while ‘Simple’ refers to the rest of the MCQs that are relatively easier.
Appendix D Human Annotation
As there is potential risk for the LLM-based Knowledge Preserving evaluation in TrinEval procedure (line 4 and line 10 in 1) that the API-based LLMs might still be able to answer the questions without sufficient knowledge since this is still prompting LLMs who have been trained on these datasets and “know” the original content, human annotation is also applied. The annotation of each MCQ encompasses three subtasks: (1) answering the question in the re-organized form, (2) answering the question in the original form, and (3) verdicting if the reformulation is Knowledge Preserving or not.
For efficient annotation, we implemented a stratified sampling procedure by selecting one MCQ per subject from all 56 MMLU subjects (as a temporary compromise for limited time, which will be expanded later) under both Qwen and GPT reorganization paradigms. This yielded 112 representative questions (2 systems × 56 subjects) for evaluation. Three human annotators independently performed dual-form assessments through: (1) Direct question answering with the reformed format first and the original format; (2) Knowledge Preservation (K.P.) scoring across two dimensions: i. Knowledge adequacy (sufficiency for accurate response), ii. removal of redundant content using a 5-point scale (1=unsatisfactory; 2=major information are missed or unnecessary information is incorporated, but part is still acceptable, 3=need to take some time to understand, but can still solve the MCQ, 4=an element properly belonging to one triplet component appears in another, but does not impact MCQ solving; 5=optimal). We use a continuous rather than binary metric to mitigate the cognitive difference of the threshold between the annotators. Inter-rater reliability was ensured through consensus-building discussions prior to formal annotation. Final scores were aggregated using mean values to further mitigate individual annotator bias. The results are shown in Tab. 4 below.
Model TrinEval correct only both correct original correct only K.P. score
Qwen 0.0% 96.296% 3.704% 4.101
GPT 0.0% 96.667% 3.333% 4.369
Table 4: Result of Human Annotation on LLM-based knowledge preserving (K.P.) evaluation in TrinEval
Notably, our analysis reveals that over 95% of correctly answered MCQs maintained consistency across both original and paraphrased formats. Furthermore, human annotators rated our paraphrased questions mean K.P. scores exceeding 4.0 (on a 5-point scale), which means that the reformulated MCQs only somehow influence the readability of humans but do not impact the solvability of the original format. This provides empirical validation that our proposed TrinEval methodology effectively preserves necessary knowledge elements from original MCQ formulations, while the influence of the LLM memorization during the evaluation is rather limited.
Experimental results under human annotation also reveal that Qwen underperforms GPT in key metrics, particularly in processing long-context texts where it occasionally omits background information (evidenced by excessive "N/A" assignments in the Context fields). This capability gap is further reflected in MCQ annotations: there are only MCQs that are merely correctly answered with the original format except for the correct MCQs with both formats.
Appendix E Elaboration on dataset-related information
We suppose that unintentional data contamination arises from crawling dataset pages (e.g., Hugging Face) during the compilation of LLM pretraining datasets. When researchers are organizing the pretraining corpus, one or more neighboring original data samples would be truncated and concatenated sequentially into a pretraining sample. Thus, according to Carlini’s theory Carlini et al. (2022a) and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
Besides, the previous study Bordt et al. (2024) also applies a similar method, the “Header Completion Test”, for tabular data memorization detection. By offering the heading rows within a CSV file, they also find that providing the preceding data samples can help to detect the memorized dataset by LLMs, which also practically proves that offering samples within the same dataset would affect the memorization evocation.
Following this, similar to the dataset name, we take the samples within the same dataset as the few-shot prompt in order to find out if the TrinEval re-organization method can avoid such memorization evocation phenomenon. Still, in Fig. 4, we can see that the blue curves remain below the green ones for the “def-fsp” setting, which proves that TrinEval can restrain the memorization evocation.
Appendix F Embedding distance of memorized and non-memorized MCQs
As there are 57 different subjects within the MMLU dataset, we believe unrelated sequences would lead to increased embedding distance even though they are among the mastered knowledge points. Here, we try to filter out the unrelated samples for each sample and thus filter the closest samples at the $21e-$ 2 level in order to make sure there are not too many unrelated samples incorporated.
To highlight this result more prominently, we employed a relatively stringent data filtering strategy in the paper and made it 1% of the closest samples in Section 4.4. In the following version, we will add this clarification part in the paper. Here in order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5%. The results are shown in Tab. 5 (RM stands for rote memorization, and GCL stands for genuine capability learning).
{tblr}
cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, vline2-3,5 = 1, vline2,4,6 = 2-8, hline1-3,9 = -, Threshold & 1% 3% 5%
Subset RM GCL RM GCL RM GCL
Llama-Qwen 29.538 16.501 32.718 17.281 33.790 18.060
Llama-GPT 28.925 17.777 31.327 18.275 32.315 19.068
Mistral-Qwen 78.663 47.648 86.003 52.995 89.598 55.525
Mistral-GPT 82.932 56.375 90.597 62.112 94.384 64.820
Vicuna-Qwen 38.761 21.006 41.396 23.189 42.490 24.185
Vicuna-GPT 37.037 22.996 41.396 23.189 42.490 24.185
Table 5: Result of averaged embedding distance of the closest MCQs under different threshold
We can see that, as we said above, the more samples we incorporated, the higher the average distance of the closest embeddings grows. Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings. This proves that the reported results are robust and solid.
Appendix G Use of AI assistants
ChatGPT [^4] https://chat.openai.com/ and Qwen [^5] https://tongyi.aliyun.com/qianwen/ were used purely for the language refinement and polishment during the paper writing process. Any content generated with the LLMs was thoroughly reviewed and approved by the authors. No new content suggested by the AI assistants was used in the paper except the original expression from the authors.

### v4

@@
-Large language models are potential rote learners
-Large language models are potential rote learners
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
-Disentangling Memorization and Capability Learning in Large Language Models
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
-Large Language Models Are Potentially Rote Memorizer
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
-Large Language Models Are Both Rote Memorizer and Capacity Learner
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
-Disentangling Rote Memorization and Capability Learning for
-Large Language Model Evaluation
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
 Large Language Models Could Be Rote Learners
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6
-1College of Computer Science and Technology, Zhejiang University
-2School of Data Science and Engineering, East China Normal University
-3State Key Laboratory of Transvascular Implantation Devices,
-The Second Affiliated Hospital Zhejiang University School of Medicine
-4School of Public Health, Zhejiang University
-5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence
-6Alibaba Cloud Computing
-xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
 Abstract
@@
 First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning.
-To disentangle them, we propose TrinEval, a novel evaluation framework that reformulates MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment.
+To disentangle them, we propose TrinEval, a novel evaluation framework reformulating MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment.
 Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
-Large Language Models Could Be Rote Learners
-Yuyang Xu1,3,5, Renjun Hu2, Haochao Ying3,4,5,🖂, Jian Wu3,4,5, Xing Shi6, Wei Lin6 1College of Computer Science and Technology, Zhejiang University 2School of Data Science and Engineering, East China Normal University 3State Key Laboratory of Transvascular Implantation Devices, The Second Affiliated Hospital Zhejiang University School of Medicine 4School of Public Health, Zhejiang University 5Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence 6Alibaba Cloud Computing xuyuyang@zju.edu.cnhaochaoying@zju.edu.cn
 footnotetext: ${}^{\text{\Letter}}$orresponding Author: Haochao Ying.
 1 Introduction
+Figure 1: MCQ-based evaluation. We observe that LLMs tend to underperform on memorized MCQs.
 The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks (Ouyang et al., 2022; OpenAI, 2024; Touvron et al., 2023).
@@
 (2023) discovers that smaller models with deliberate pre-exposure could outperform their larger counterparts, thereby contradicting widely accepted scaling laws.
-Figure 1: MCQ-based LLM evaluation. We observe that LLMs tend to underperform on memorized MCQs.
 To mitigate the issue, Zhou et al.
@@
 3.1 Pre-investigation of LLM Capability w.r.t. Memorization
+Figure 2: Model performance on memorized and non-memorized subsets of MMLU, where ‘0s’ and ‘5s’ stand for zero- and five-shot prompting, respectively.
 Benchmark contamination often leads to inflated performance estimate.
@@
 The detailed statistics of questions across subsets are given in Table 1 of Appendix A, and we also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO (Wang et al., 2024).
-Figure 2: Model performance on memorized and non-memorized subsets of MMLU, where ‘0s’ and ‘5s’ stand for zero- and five-shot prompting, respectively.
 We then compute the accuracy (ACC) of tested LLMs by subsets as a proxy of model performance under different memorization conditions.
@@
 In this section, we conduct extensive experiments to answer the following questions:
-Q1. Is TrinEval knowledge-preserving in order to fulfill knowledge assessment?
-Q2. Can TrinEval reduce memorization effects during capability evaluation?
-Q3. What does TrinEval reveal about LLMs’ rote memorization and genuine capability?
+Q1. Does TrinEval effectively preserve knowledge to enable accurate knowledge assessment?
+Q2. To what extent does TrinEval mitigate memorization effects during capability evaluation?
+Q3. What insights does TrinEval provide into the distinction between rote memorization and genuine capability in LLMs?
 4.1 Experiment Setup
 Models.
-We utilize API-based commercial LLMs, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024) for question reformulation by TrinEval.
-Model evaluation is conducted on open-source LLMs due to limited budgets, and we experiment with three popular LLMs including Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b).
-All the three LLMs are accessed from Huggingface and implemented with transformers library, we thus could obtain the log-probability of output token for fine-grained study.
-Throughout our tests, we use the default generation parameters and adopt greedy decoding to enhance reproducibility.
+We employ API-accessible commercial LLMs for question reformulation in TrinEval, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024).
+For model evaluation, we utilize open-source LLMs to obtain token-level logits.
+Our experiments focus on three widely adopted models: Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b).
+These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis.
+All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory.
+Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
 Benchmarks.
-We evaluate LLM on the widely used MMLU (Hendrycks et al., 2020) benchmark.
-MMLU consists of 57 subjects from areas including STEM, humanities, social sciences, and others, enabling comprehensive evaluation of LLM capacity.
-As there are duplicated MCQs across different subjects, we eliminate them and obtain 14,006 MCQs as the test set.
+We evaluate LLMs on the widely used MMLU benchmark (Hendrycks et al., 2020), in light of growing evidence of data contamination (Touvron et al., 2023) [^3] https://hitz-zentroa.github.io/lm-contamination/ in many recent LLMs on this data.
+MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities.
+We find duplicated MCQs across subjects on MMLU, and thus de-duplicate the dataset, resulting in a test set of 14,006 unique questions.
 Evaluation.
-With commercial LLMs, we evaluate model performance by extracting answers with regular expressions.
-For open-source LLMs, we access the output probability of the first generated token (e.g., option IDs A/B/C/D) to obtain a quantitative performance result.
+With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions.
+For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metrics.
 4.2 Q1. Is TrinEval Knowledge-preserving?
-We first verify whether the reformulation by TrinEval is knowledge-preserving in order to fulfill knowledge assessment.
-To achieve this, our primary objective is to validate that the reformulation approach (1) does not lose key information that results in previous correctly-answered questions being answered incorrectly and (2) does not introduce anomalous or unexpected information that results in inflated performance.
-Figure 3: Knowledge-preserving validation for reformulation by TrinEval.
-4,343 and 4,645 qualified MCQs are obtained with Qwen and GPT after reformulation.
-We test Qwen and GPT in these qualified subsets.
-The green and blue circles stand for the correctly answered MCQs in TrinEval and original formats, respectively.
-(a) Evocation with Llama2 and dev-fsp.
-(b) Evocation with Mistral and dev-fsp.
-(c) Evocation with Vicuan and dev-fsp.
-(d) Evocation with Llama2 and seq-fsp.
-(e) Evocation with Mistral and seq-fsp.
-(f) Evocation with Vicuan and seq-fsp.
-Figure 4: The results of memorization evocation under various dataset-related information context, with blue and green curves referring to the memorization difference $\Delta F_{m}$ in the TrinEval and original formats, respectively.
-In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the training set, and with few-shot prompt from the test set ahead of the current testing question.
-These curves indicate the growing memorization metric $\Delta F_{m}$ with the stronger dataset-related information in general.
-However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context are consistently lower than those in the original format, e.g., ‘clean’.
-Upon completing the complete TrinEval reformulation process, we ultimately obtained 4,343 MCQs and their corresponding knowledge entities, attributes, and contexts that met our criteria using Qwen, as well as 4,645 qualified MCQs and their respective triplets using GPT.
-We then instruct Qwen and GPT to answer these respective questions in both the original (baseline) and restated triplet form.
-The prompts and an MCQ example are available in Table 2 in Appendix and the results are shown in Fig.
+Figure 3: Knowledge-preserving validation of TrinEval reformulation.
+The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively.
+Best viewed in color.
+We begin by verify whether TrinEval’s reformulation preserves essential knowledge, a prerequisite for reliable knowledge assessment.
+Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance.
+Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding knowledge entities, attributes, and contexts using Qwen, and 4,645 MCQs with associated triplets using GPT.
+We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats.
+Example MCQs and the questioning prompt templates in both formats are provided in Table 2 in the Appendix; results are summarized in Fig.
 3.
-We can observe that for Qwen, 92.95% of correctly answered MCQs in the TrinEval format maintain their accuracy in the original format, while for GPT, 90.05% of qualified MCQs are answered correctly, with 95% of these maintaining accuracy in the original format.
-That is, for both Qwen and GPT, we can infer that the correctly answered MCQs from the qualified ones in TrinEval format can be regarded as a subset of the correctly answered MCQs with the original MCQ format.
-This proves that the proposed TrinEval reformulation method does not incorporate extra information that leads to additional capability of LLMs.
-On the other hand, the intersection MCQs between the correctly answered in two formats also make up of around 95% of the MCQs correctly answered in the original MCQ format, which proves that the TrinEval incorporates all the necessary information to answer the question.
-In conclusion, our TrinEval effectively retains the LLMs’ problem-solving capability compared to the original MCQ text.
+Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats, respectively.
+Notably, the majority of these are answered correctly in both the original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats.
+Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting that the TrinEval reformulation does not omit essential information.
+Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone.
+This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance.
+Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for accurate answering, satisfying the requirements for reliable capability evaluation.
 4.3 Q2. Can TrinEval Reduce Memorization?
-In this subsection, we aim to validate whether the proposed TrinEval can eliminate the unnecessary memorization of LLMs, and thus demonstrate enhanced robustness against various perturbations.
-To answer this question, following Bordt et al.
-(2024), we deliberately incorporate the dataset-related information into the context and evaluate whether the TrinEval reformulation can suppress the growing memorization level with memorization evocation of different extent and can reveal the genuine capabilities of LLMs.
-We incorporate the dataset-related information into the context, i.e., the name of the dataset, and the few-shot prompt of samples within the same dataset for the memorization-evocation perturbation (see Appendix E).
-Here we use Llama, Mistral, and Vicuna as the tested LLMs since we access the output probabilities to compute the memorization metric $F_{m}$.
-As there is no specific zero point of $F_{m}$ indicating the absolute-no memorization of MCQs given an LLM, in order to better visualize the difference between the proposed TrinEval and the original MCQ baseline format, we take the $F_{m}$ with vanilla MCQ (i.e., original MCQ format without any dataset-related prompt) as the baseline and visualize the averaged difference between the $F_{m}$ of the tested format and the baseline.
-(a) $F_{m}$ v.s. $F_{c}$ with Llama2 based on Qwen-extracted triplets.
-(b) $F_{m}$ v.s. $F_{c}$ with Mistral based on Qwen-extracted triplets.
-(c) $F_{m}$ v.s. $F_{c}$ with Vicuna based on Qwen-extracted triplets.
-(d) $F_{m}$ v.s. $F_{c}$ with Llama2 based on GPT-extracted triplets.
-(e) $F_{m}$ v.s. $F_{c}$ with Mistral based on GPT-extracted triplets.
-(f) $F_{m}$ v.s. $F_{c}$ with Vicuna based on GPT-extracted triplets.
-Figure 5: The distribution of MCQs based on Memorization metric $F_{m}$ v.s.
-the Capability metric $F_{c}$.
-According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs equally into five groups and visualize the distribution of MCQs with the heatmap from weak to strong.
-Specifically, for the memorization-evocation permutation, we progressively enhance the prompt context for memory evocation, starting from merely providing the dataset name, to offering samples within the same dataset as few-shot prompts (including the training set of the dataset-dev, and the preceding samples adjacent to the test sample-seq), and finally to providing both as context.
-Here for each tested MCQ, we calculate the difference between the $F_{m}$ given the corresponding memorization evocation prompt and the $F_{m}$ with the vanilla MCQ baseline.
-Fig.
-4 shows the curve based on the average difference of each MCQ.
-From this figure, as stated by Bordt et al.
-(2024), we can see that $F_{m}$ is growing with the stronger dataset-related context.
-When providing more specific context related with the test dataset, the LLMs tend to exhibit stronger memorization of the MCQs.
-Specially, for all three open-source LLMs, the $\Delta F_{m}$ curve of TrinEval is below the curve of the original MCQ baseline.
-More importantly, the $F_{m}$ of TrinEval with the strongest memorization evocation is still below the vanilla MCQ baseline, which proves that TrinEval can effectively eliminate the memorization from LLMs.
-4.4 Q3. TrinEval’s Findings on Memorization and Capability
-Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings.
-‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$.
-Further results are shown in Appendix F
-In this subsection, we aim to explicitly study the relationship between the memorization and the capability of LLMs with the metrics $F_{m}$ and $F_{c}$.
-As the commercial-API-based LLMs do not provide the output probability of the whole vocabulary, we mainly use the open-source LLMs to compute these two metrics.
-After obtaining the $F_{m}$ and $F_{c}$ of each MCQ, we separate all the qualified MCQs into 5 equal groups.
-Finally, we utilize the heatmap to reveal the relationship between the capability and the memorization of the tested LLMs.
-As shown in Fig.
-5, most of the MCQs concentrate on the lower left corner and the upper right corner of the heatmap.
-Specifically, for $F_{m}$ v.s.
-$F_{c}$ with Llama2 based on Qwen-extracted triplets, MCQs within the lower left $2\times 2$ squares and the upper right $2\times 2$ squares make up of the 38.57% of all the tested MCQs with a Pearson-correlation of -0.7755 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 74.17% of all the tested MCQs with a Pearson-correlation of -0.8124 (p-value $<0.05$).
-For the results with Mistral based on Qwen-extracted triplets, MCQs within the lower left and upper right $2\times 2$ squares make up of the 44.90% of all the tested MCQs with a Pearson-correlation of -0.8722 (p-value $<0.05$), while the MCQs within the lower left and upper right $3\times 3$ squares make up of the 80.82% of all the tested MCQs with a Pearson-correlation of -0.8794 (p-value $<0.05$).
-More results are shown in Tab.
-3.
-This evidence indicates that MCQs with lower memorization levels tend to exhibit better problem-solving capabilities of LLMs, while those with higher memorization levels are associated with reduced performance in solving tasks.
-Next, we hypothesize that the LLMs are potential rote learners through the human memory system, which has been characterized by two fundamental components: Long-Term Memorization (LTM) and Short-Term Memorization (STM) Shiffrin (2003).
+(a) Evoc. with Llama2 and dev-fsp.
+(b) Evoc. with Mistral and dev-fsp.
+(c) Evoc. with Vicuna and dev-fsp.
+(d) Evoc. with Llama2 and seq-fsp.
+(e) Evoc. with Mistral and seq-fsp.
+(f) Evoc. with Vicuna and seq-fsp.
+Figure 4: The results of memorization evocation (evoc.) under various dataset-related context, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively.
+In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question.
+These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general.
+However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
+In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization in LLMs, thereby isolating and revealing the models’ genuine capabilities under various circumstances.
+Following the methodology of Bordt et al.
+(2024), we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
+To evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset (see Appendix E).
+For this set of experiments, we focus on LLaMA, Mistral, and Vicuna, as their open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$.
+Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we use the $F_{m}$ value obtained from the original MCQ format without any additional context (i.e.,, the vanilla MCQ) as a reference baseline.
+We then visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
+Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction.
+Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al.
+(2022a), we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization.
+Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context.
+For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats.
+Results are illustrated in Fig.
+4.
+Consistent with the observations by Bordt et al.
+(2024), the results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects.
+However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context used.
+Notably, even under the strongest memorization-evoking setting, TrinEval’s absolute $F_{m}$ still stays below the vanilla baseline.
+These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling spurious recall from genuine model capability.
+4.4 Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability
+(a) $F_{m}$ v.s. $F_{c}$ with Llama on Qwen-extracted triplets.
+(b) $F_{m}$ v.s. $F_{c}$ with Mistral on Qwen-extracted triplets.
+(c) $F_{m}$ v.s. $F_{c}$ with Vicuna on Qwen-extracted triplets.
+(d) $F_{m}$ v.s. $F_{c}$ with Llama on GPT-extracted triplets.
+(e) $F_{m}$ v.s. $F_{c}$ with Mistral on GPT-extracted triplets.
+(f) $F_{m}$ v.s. $F_{c}$ with Vicuna on GPT-extracted triplets.
+Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs.
+capability metric $F_{c}$.
+According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs into 25 groups and visualize the MCQ distribution from weak to strong with heatmaps.
+In this subsection, we aim to explicitly examine the relationship between memorization and genuine problem-solving capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability).
+Since commercial API-based models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics.
+After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution.
+We then use heatmaps to visualize the distribution of samples across these groups, thereby revealing the relationship between memorization and capability for each tested model.
+As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap.
+This pattern is evident and consistent across our tests.
+For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value $<0.05$) for $F_{m}$ vs.
+$F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value $<0.05$).
+Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value $<0.05$), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value $<0.05$).
+We take the MCQs within the upper-right 2 X 2 squares in Fig.
+5 as the ones memorized by rote and compute the ratio of 19.6% by averaging the percentage of all evaluated LLMs.
+Similarly, 20.5% is computed as the ratio of MCQs within the lower-left 2 X 3 squares with genuine capability.
+Additional statistics results are reported in Table 3 in Appendix.
+These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall.
+To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) (Shiffrin, 2003).
 Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability.
@@
 This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts.
-As illustrated in Allen-Zhu and Li (2023) and Ovadia et al.
-(2023), LLMs trained with multiple rephrased corpus tend to perform better than LLMs trained with only the original corpus.
-When providing only one format of training corpus, similar to the STM system, LLMs tend to memorize the corpus at token-level rather than knowledge-level.
-In other words, LLMs encode these corpora at a shallow level with the original format.
-After questions are rephrased with methods like our proposed TrinEval, the input corpus seems connected with the known knowledge like the LTM for structured storage and enables sophisticated information generalization.
-We show more detailed results in Appendix C.
-To further validate our hypothesis, we compute the embeddings of MCQs within the qualified MMLU dataset and average the distance between the other closest $1\%$ MCQs.
-We visualize the mean distance of MCQs within the lower left and upper right $2\times 2$ squares in Fig.
+This dichotomy aligns with recent observations in LLMs.
+As shown by Allen-Zhu and Li (2023) and Ovadia et al.
+(2023), models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations.
+When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge.
+In this sense, LLMs are potential rote learners.
+Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation.
+Further experimental details and results are provided in Appendix C.
+Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings.
+‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$.
+Further results are shown in Appendix F.
+To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% of samples.
+Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus.
+As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs.
+By selecting only the top 1% nearest neighbors, we aim to retain the samples that are semantically-correlated in knowledge description.
+We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions of the heatmap in Fig.
 5.
-The results are shown in Fig.
+The corresponding results are presented in Fig.
 6.
-We surprisingly find that the averaged distance of the Genuine Capability Learning MCQs (i.e., MCQs within the upper right $2\times 2$ squares) is almost half as much as the distance of the Rote Memorization MCQs (i.e., MCQs within the lower left $2\times 2$ squares).
-The result hints that the memorized MCQs are sparsely encoded by MCQs while the non-memorized ones share common embeddings, which is again coincident with the findings of the STM and LTM.
+Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs.
+This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters.
+This observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations.
+The results of alternative thresholds are shown in Table 5.
 Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.
@@
 1.
-Prompt template for extraction:
-You are an expert of multiple choice questions of MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A.
-Prompt template for pre-investigation on LLM Memorization w.r.t. Capability:
+Prompt template for extraction: Prompt template for pre-investigation on LLM Memorization w.r.t. Capability:
+You are an expert of multiple choice questions of MMLU dataset.
+The following are multiple-choice questions (with answers) about [subject].
+[examples] [question] Options: A.
 You are an expert of multiple choice questions of MMLU dataset.
@@
 **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
-Original MCQ TrinEval MCQ
+Original MCQ Original MCQ Example
 You are an expert on multiple choice questions of [subject].
@@
 [option C] D.
-[option D] You are an expert on multiple choice questions of [subject].
+[option D] You are an expert on multiple choice questions of high school computer science.
+Analyze the given question and the given options.
+Determine the correct answer option to the question.
+Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options.
+Generate the option ID of the correct option (answer).
+- **Question:** Which of the following is usually NOT represented in a subroutine’s activation record frame for a stack-based programming language?
+- **Options:** A.
+Values of local variables B.
+A heap area C.
+The return address D.
+Stack pointer for the calling activation record
+TrinEval MCQ TrinEval MCQ Example
+You are an expert on multiple choice questions of [subject].
 Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question.
@@
 [option C] D.
-[option D]
-Original MCQ Example TrinEval MCQ Example
-You are an expert on multiple choice questions of high school computer science.
-Analyze the given question and the given options.
-Determine the correct answer option to the question.
-Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options.
-Generate the option ID of the correct option (answer).
-- **Question:** Which of the following is usually NOT represented in a subroutine’s activation record frame for a stack-based programming language?
-- **Options:** A.
-Values of local variables B.
-A heap area C.
-The return address D.
-Stack pointer for the calling activation record You are an expert on multiple choice questions of high school computer science.
+[option D] You are an expert on multiple choice questions of high school computer science.
 Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question.
@@
 3.
+All person correlation values are computed with p-values at the 0.05 level.
 (a) Probability and p-value with Llama2 based on GPT.
@@
 Thus, according to Carlini’s theory Carlini et al.
-(2022a) and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
+[2022a] and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
 Besides, the previous study Bordt et al.
-(2024) also applies a similar method, the “Header Completion Test”, for tabular data memorization detection.
+[2024] also applies a similar method, the “Header Completion Test”, for tabular data memorization detection.
 By offering the heading rows within a CSV file, they also find that providing the preceding data samples can help to detect the memorized dataset by LLMs, which also practically proves that offering samples within the same dataset would affect the memorization evocation.
@@
 This proves that the reported results are robust and solid.
-Appendix G Use of AI assistants
-ChatGPT [^4] https://chat.openai.com/ and Qwen [^5] https://tongyi.aliyun.com/qianwen/ were used purely for the language refinement and polishment during the paper writing process.
-Any content generated with the LLMs was thoroughly reviewed and approved by the authors.
-No new content suggested by the AI assistants was used in the paper except the original expression from the authors.

### v5

@@
 Abstract
-Multiple-choice question (MCQ) benchmarks are widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination.
-In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle genuine capability acquisition from superficial memorization in LLM evaluation.
-First, by analyzing model performance under different memorization conditions, we uncover a counterintuitive trend: LLMs perform worse on memorized MCQs than on non-memorized ones, indicating the coexistence of two distinct learning phenomena, i.e., rote memorization and genuine capability learning.
-To disentangle them, we propose TrinEval, a novel evaluation framework reformulating MCQs into an alternative trinity format, reducing memorization while preserving knowledge assessment.
-Experiments validate TrinEval’s effectiveness in reformulation, and its evaluation reveals that common LLMs may memorize by rote 20.5% of knowledge points (in MMLU on average).
-footnotetext: ${}^{\text{\Letter}}$orresponding Author: Haochao Ying.
-1 Introduction
-Figure 1: MCQ-based evaluation. We observe that LLMs tend to underperform on memorized MCQs.
-The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks (Ouyang et al., 2022; OpenAI, 2024; Touvron et al., 2023).
-As LLMs continue to improve, evaluating their genuine capacities has emerged as a fundamental challenge, necessitating proper methodologies to ensure fairness and robustness (Ganguli et al., 2023; Liu et al., 2023b).
-Among the developed methods, multiple-choice question (MCQ) benchmarks have become a standard approach for evaluation.
-Typically, LLMs are presented with a question and a fixed set of answer choices, requiring them to select the most appropriate option (see Fig.
+Benchmark-based evaluation, e.g., multiple-choice questions (MCQs) and open-ended questions (OEQs), is widely used for evaluating Large Language Models (LLMs), yet their reliability is undermined by benchmark contamination.
+When pre-exposed to testing benchmarks during training, less capable LLMs have been found to achieve inflated performance, thereby yielding erroneous results.
+In this study, we reframe contamination as an inherent aspect of learning and seek to disentangle and expose genuine capability acquisition from superficial memorization in LLM evaluation.
+Following this, firstly, by analyzing model performance under different memorization conditions of MCQs, we uncover a counterintuitive trend: LLMs perform worse on memorized benchmarks than on non-memorized ones, indicating the coexistence of two learning phenomena, i.e., rote memorization and genuine capability learning.
+To disentangle them, we propose TrinEval, a novel evaluation framework that reformulates MCQs into an alternative knowledge-centric trinity format, reducing memorization while preserving inherent knowledge, enabling the evaluation of genuine capability in the presence of memorization.
+Extensive experiments validate the effectiveness and robustness of TrinEval in reformulating benchmarks, and the evaluation results further reveal that mainstream LLMs rely on rote memorization for an average of 19.6% of knowledge points across the MMLU and the GSM8K dataset.
+I Introduction
+The rapid advancement of Large Language Models (LLMs), driven primarily by large-scale pre-training on massive datasets, has endowed these models with remarkable proficiency across diverse tasks [47, 13, 7].
+As LLMs continue to improve, evaluating their genuine capacities has emerged as one of the fundamental challenges, necessitating proper methodologies to ensure fairness and robustness [48, 49].
+Among the developed methods, benchmark-based evaluation has become a standard approach to evaluate LLMs.
+Typically, taking multiple-choice questions (MCQs) as an example, LLMs are presented with a question and a fixed set of optional choices, requiring them to select the most appropriate answer (see Fig.
 1 for illustration).
 This format enables straightforward performance measurement through accuracy metrics and could cover a wide range of subjects.
-However, despite their widespread adoption, MCQ-based evaluation raises concerns about reliability due to benchmark contamination (Li and Flanigan, 2024; Kim et al., 2024), i.e., test data unintentionally appears in training corpora and models may exploit memorized content rather than demonstrating genuine understanding, inflating their apparent capabilities.
+However, despite their widespread adoption, benchmark-based evaluation raises concerns about reliability due to benchmark contamination [35, 36], i.e., test data unintentionally appears in training corpora, and LLMs may exploit memorized content rather than demonstrating genuine understanding, inflating their apparent capabilities.
 For instance, Zhou et al.
-(2023) discovers that smaller models with deliberate pre-exposure could outperform their larger counterparts, thereby contradicting widely accepted scaling laws.
-To mitigate the issue, Zhou et al.
-(2023) advocates the removal of benchmark datasets from pre-training corpora.
+[34] discovers that smaller models with deliberate pre-exposure could outperform larger counterparts, thereby contradicting widely accepted scaling laws.
+Figure 1: Benchmark-based evaluation. We observe that LLMs tend to underperform on memorized questions.
+To mitigate this issue, Zhou et al.
+[34] advocates the removal of benchmarks from pre-training corpora.
 However, this strategy conflicts with the fundamental objective of large-scale pre-training, which aims to maximize model performance by exposing LLMs to as much data as possible.
 From a broader perspective, human learning also involves problem-solving through practicing on similar questions, e.g., exam preparation.
-While rote memorization of specific questions and answers merely lead to short-term success, repeated practicing can also facilitate deeper conceptual understanding.
-Inspired, rather than viewing benchmark contamination as a flaw to be eradicated, which is a nearly impossible task at scale (Sainz et al., 2023; Bordt et al., 2024), we argue that it is an inherent aspect of learning and should be accounted for in evaluation.
-Therefore, this study shifts its focus to evaluating LLMs in the presence of contamination, aiming to distinguish genuine capability gains from superficial memorization effects.
-The explicit disentangling of these two learning effects remains largely unexplored in MCQ-based evaluation, yet we believe it marks a crucial step towards developing more rigorous and unbiased evaluation methodologies.
-To investigate the effects of superficial memorization in LLM evaluation, we compare model performance under different memorization conditions.
-Inspired by membership inference attacks (MIA) (Carlini et al., 2022a, 2021), we define superficial memorization as an LLM’s ability to verbatim reproduce content, e.g., MCQs in our case.
-Using this criterion, we partition the MMLU benchmark (Hendrycks et al., 2020) [^1] Selected for its popularity and documented data contamination in widely used LLMs (Sainz et al., 2023).
-into memorized and non-memorized subsets and evaluate three open-source LLMs [^2] Llama2-7B (Touvron et al., 2023), Mistral-7B-v0.2 (Jiang et al., 2023) and Vicuna-v1.5-7B (Zheng et al., 2023b).
-on both.
+While rote memorization of specific questions and answers merely leads to short-term success, repeated practice can also facilitate deeper conceptual understanding.
+Thus, rather than viewing benchmark contamination as a flaw to be eradicated, which is a nearly impossible task at scale [3, 4], we argue that it is an inherent aspect of learning and should be accounted for in evaluation.
+Therefore, this study shifts focus to evaluating LLMs in the presence of contamination, aiming to distinguish genuine capability gains from superficial memorization effects.
+The explicit disentangling of these two learning effects remains largely unexplored in benchmark-based (MCQ-based, as an example) evaluation, yet we believe it marks a crucial step towards developing more rigorous and unbiased evaluation methodologies.
+To investigate superficial memorization effects in LLM evaluation, we compare model performance under different memorization conditions.
+Inspired by membership inference attacks (MIA) [1, 37], we consider the intuitive manifestation of superficial memorization is the LLM’s ability to verbatim reproduce content, e.g., extracting options of MCQs in our case.
+With the criterion, we partition the MMLU benchmark [6] [^1] Selected for its popularity and documented data contamination in widely used LLMs [3].
+into memorized and non-memorized subsets and evaluate three open-source LLMs [^2] Llama2-7B [7], Mistral-7B-v0.2 [11] and Vicuna-v1.5-7B [12]..
 Surprisingly, results reveal a consistent yet counterintuitive trend: LLMs perform worse on memorized MCQs than on those not (see Fig.
@@
 2 for results).
-This challenges the assumption that memorization improves model performance and suggests the coexistence of two distinct learning phenomena in LLMs: rote memorization, where models recall content verbatim without true understanding, and genuine capability learning, where they internalize underlying knowledge.
-The preliminary investigation has several limitations.
-First, the binary classification of MCQs as either memorized or non-memorized oversimplifies the nuances of memorization, potentially overlooking intermediate cases.
-Second, we rely on accuracy to measure performance, which is inherently unreliable.
-Third, our analysis could not reveal the mutual effects between rote memorization and capability learning.
-To address these challenges, we propose TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization.
-TrinEval employs a query-based probing (q-probing) mechanism Allen-Zhu and Li (2023) that reformulates MCQs into an alternative trinity format, i.e., entity-attribute-context.
-This could prevent direct content recall while preserving knowledge assessment.
+This challenges the assumption that memorization inflates model performance and suggests the coexistence of two learning phenomena in LLMs: rote memorization, where models recall content verbatim without true understanding, and genuine capability learning, where they internalize underlying knowledge.
+Still, our preliminary investigation needs further exploration.
+First, binary classification of benchmarks as either memorized/non-memorized and capable/incapable oversimplifies the nuances of memorization and capability, potentially overlooking intermediate cases.
+Second, our analysis could not reveal the numerical dependency between rote memorization and capability learning.
+To address these challenges, we propose TrinEval, a novel evaluation framework designed to provide a reliable measure of LLM performance by minimizing the influence of rote memorization.
+TrinEval employs a query-based probing (q-probing) mechanism [18] that reformulates benchmarks into an alternative trinity format, i.e., entity-attribute-context, which prevent direct content recall while preserving knowledge assessment, and exposes the genuine capability of LLMs.
 Through experiments, we demonstrate that TrinEval’s reformulation is knowledge-preserving, i.e., maintaining testing problems’ inherent knowledge requirements without introducing extra cues, and could effectively reduce memorization.
-Combined with a continuous superficial memorization quantification metric, TrinEval reveals the in-robustness of LLMs’ capability learning, e.g., with MMLU, tested open-sourced LLMs only mastered 19.6% of knowledge points while 20.5% are memorized by rote in the meanwhile, shedding light on the necessity for further optimization.
-2 Related Work
-2.1 LLM Evaluation on MCQ Benchmarks
-The rapid advancement of LLMs has driven their expansion into diverse domains, necessitating robust and fair evaluation methodologies (Zheng et al., 2023b; Hu et al., 2025) and platforms (Contributors, 2023; Chiang et al., 2024).
-Among these, evaluating on MCQ benchmarks emerges as a widely adopted approach due to the ease of validation and standardized comparison across models (Hendrycks et al., 2020; Wang et al., 2024; Zhong et al., 2023; Huang et al., 2024).
+Combined with a continuous superficial memorization quantification metric, TrinEval reveals the in-robustness of LLMs’ capability in benchmark evaluation.
+With MMLU, tested open-sourced LLMs only mastered 19.0% of knowledge points while 19.6% are memorized by rote in the meanwhile, while for OEQs in GSM8K, LLMs only mastered 17.12% of knowledge points while 17.30% are rote memorized, shedding light on the necessity for further optimization on the LLMs development.
+Our main contributions are as follows:
+  • We propose a novel evaluation paradigm that assesses the genuine capability of LLMs under the default condition of benchmark contamination.
+  • We introduce TrinEval, a framework that gets access to the genuine knowledge stored within tokens through textual reformulation and explores LLMs’ true understanding based on verbal queries.
+  • Through TrinEval’s memorization-capability decoupling and quantitative analysis, we reveal a counterintuitive but general finding: the more LLMs memorize, the worse their actual problem-solving capability becomes.
+II Related Work
+II-A LLM Evaluation on Benchmarks
+The rapid advancement of LLMs has driven their expansion into diverse domains, necessitating robust and fair evaluation methodologies [12, 8] and platforms [24, 25].
+Among these, evaluating on MCQ benchmarks emerges as a widely adopted approach due to the ease of validation and standardized comparison across models, including MMLU [6], a large-scale multi-task benchmark covering 57 subjects spanning STEM, humanities, social sciences, and professional disciplines, designed to assess general knowledge and reasoning ability; AGIEval [27], a human-centric benchmark constructed from real-world public examinations such as college entrance exams and certification tests, focusing on practical knowledge application and high-level cognitive skills; C-Eval [28], a comprehensive Chinese multi-level evaluation suite covering middle school, high school, university, and professional domains, tailored for evaluating foundation models in Chinese understanding and reasoning; and MMLU-Pro [5], an enhanced version of MMLU that filters out simplistic questions and retains complex, model-challenging items to better measure robust high-order reasoning capacity.
 However, MCQ-based evaluations are not without limitations.
-Biases in LLM responses have been extensively studied (Dai et al., 2024), revealing issues such as social biases (Salewski et al., 2024; Liu et al., 2023a) and order sensitivity (Akter et al., 2023).
-To mitigate the latter, Pride (Zheng et al., 2023a) estimates the option positional bias after option permutation.
+Biases in LLM responses have been extensively studied [26], revealing issues such as social biases [30, 29] and order sensitivity [31].
+To mitigate the latter, PriDe [22] estimates the option positional bias after option permutation.
 To examine mastery of knowledge, Zhao et al.
-(2023) applies a hypothesis testing method and checks rephrased-context consistency for a given question.
-Benchmark contamination is arguably the most severe challenge for MCQ-based evaluations, which may result in misleadingly inflated performance (Zhou et al., 2023; Li and Flanigan, 2024).
-To address this, prior studies have explored data filtering, frequently-updated test sets (White et al., 2025), and data perturbation (Li et al., 2024).
+[32] apply a hypothesis testing method and check rephrased-context consistency for a given question.
+Benchmark contamination is arguably the most severe challenge for MCQ-based evaluations, which may result in misleadingly inflated performance [34, 35].
+To address this, prior studies have explored data filtering, frequently-updated test sets [9], and data perturbation [10], yet none of these can fundamentally resolve the issue.
+Beyond MCQs-based benchmarks, open-ended questions (OEQs) evaluation has gained increasing attention for assessing LLMs’ generative capabilities, reasoning chains, and factual consistency in free-form responses.
+Representative open-ended benchmarks include GSM8K [53], which focuses on multi-step mathematical reasoning with detailed chain-of-thought annotations; MATH [6], covering advanced algebraic, geometric, and combinatorial problems; TruthfulQA [57], designed to measure truthfulness and avoid falsehoods in generation; and CommonsenseQA 2.0 [59] for commonsense understanding in free-response settings.
+These benchmarks require LLMs to produce coherent, logically complete, and factually accurate answers rather than selecting from fixed options, making them more aligned with real-world applications.
+Still, OEQ-based evaluation faces similar challenges from benchmark contamination.
+Since test samples, including full questions, reasoning steps, and reference answers, are frequently leaked into training corpora, LLMs may directly reproduce memorized rationales instead of performing genuine reasoning [35, 3].
+Also, evaluating OEQ responses introduces additional difficulties: automated metrics such as BLEU, ROUGE, or BERTScore often fail to capture logical correctness, while human evaluation is costly, inconsistent, and difficult to scale [8, 12].
+Recent work has attempted to mitigate these issues by perturbing question templates [10], using dynamic test sets [9], or applying model-based judges [24, 8], yet few methods [60] explicitly disentangle rote memorization from genuine reasoning ability in open-ended generation and provide a commonly agreed evaluation metric.
+As a result, reliably evaluating the true problem-solving capacity of LLMs under contamination remains an open challenge for both MCQ and OEQ benchmarks.
 In this paper, instead of attempting to eliminate contamination, we evaluate LLMs under its presence, aiming to distinguish genuine capability gains from superficial memorization effects.
-This marks a new perspective of LLM evaluation, revealing the extent to which models truly understand concepts rather than merely memorizing data.
-2.2 LLM Memorization
+This marks a new perspective on LLM evaluation, revealing the extent to which models truly understand concepts rather than merely memorizing data.
+II-B LLM Memorization
 Membership inference attacks (MIA) are commonly used to determine whether a specific sample was present in a model’s training data.
 Initially studied in smaller models, Carlini et al.
-(2022b) investigates deep learning memorization mechanisms by identifying and removing easily detectable memorized samples.
-In the context of LLMs, MIA has been employed to assess privacy risks, revealing that both open- and closed-source models can leak sensitive personal data when provided with related prompts (Kim et al., 2024).
+[33] investigates deep learning memorization mechanisms by identifying and removing easily detectable memorized samples.
+In the context of LLMs, MIA has been employed to assess privacy risks, revealing that both open- and closed-source models can leak sensitive personal data when provided with related prompts [36].
 Beyond privacy concerns, Carlini et al.
-(2022a) formally defines LLM memorization as a model’s ability to verbatim generate text sequences following a prefix prompt.
-Using this definition, several studies (Sainz et al., 2023; Bordt et al., 2024; Carlini et al., 2021) have examined mainstream LLMs, confirming widespread test data leakage across popular benchmarks.
-To quantify memorization strength, researchers (Shi et al., 2023; Zhang et al., 2024; Oren et al., 2023; Carlini et al., 2019) have further explored methods such as analyzing token probability distributions in generated outputs.
+[1] formally defines LLM memorization as a model’s ability to verbatim generate text sequences following a prefix prompt.
+Using this definition, several studies [3, 4, 37] have examined mainstream LLMs, confirming widespread test data leakage across popular benchmarks.
+To quantify memorization strength, researchers [16, 17, 38, 39] have further explored methods such as analyzing token probability distributions in generated outputs.
 However, while these studies extensively analyze LLM memorization, few explicitly investigate how memorization influences an LLM’s problem-solving ability.
 In contrast, our work focuses on their interplay, presenting a more rigorous approach to fair and reliable LLM evaluation.
-3 Methodology
-3.1 Pre-investigation of LLM Capability w.r.t. Memorization
-Figure 2: Model performance on memorized and non-memorized subsets of MMLU, where ‘0s’ and ‘5s’ stand for zero- and five-shot prompting, respectively.
+III Methodology
+III-A Pre-investigation of LLM Capability w.r.t. Memorization
 Benchmark contamination often leads to inflated performance estimate.
-This phenomenon is commonly attributed to models memorizing specific questions and answers rather than demonstrating genuine problem-solving abilities.
+This phenomenon is commonly attributed to models memorizing specific question and answer sequence rather than demonstrating genuine problem-solving abilities.
 However, the extent to which and how memorization influences LLM performance remains unclear.
@@
 non-memorized subsets, we aim to reveal the role of memorization in LLM evaluation and establish a foundation for more rigorous assessment methodologies.
+In this paper, we mainly discussed the LLM Capability w.r.t.
+Memorization via MCQs as an example [^3] We also discuss and validate the extension of TrinEval on OEQs in Sec.
+IV-D..
 Formally, we define an MCQ as $x=\{x_{Q},x_{O},x_{W}\}$, where $x_{Q}$, $x_{O}$, and $x_{W}$ refer to the question, options, and ground-truth answer, respectively.
 Following the memorization definition from Carlini et al.
-(2022a), we say an MCQ $x$ is memorized by LLM $G$ if $G$ can extract/generate the content of options $x_{O}$ exactly given question $x_{Q}$.
-In practice, we incorporate meta-information (e.g., benchmark name) and 5-shot examples to recall memory and use greedy decoding (i.e., temperature fixed to 0) during extraction (Bordt et al., 2024; Sainz et al., 2023) (refer to Appendix A for the complete prompt).
-Using MMLU (Hendrycks et al., 2020) as the evaluation benchmark, we divide the test set MCQs into memorized and non-memorized subsets, where the memorized subset consists of 909–982 questions (accounting for 6.5%–7.0% of the total 14,006) depending on the tested LLMs Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B.
-The detailed statistics of questions across subsets are given in Table 1 of Appendix A, and we also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO (Wang et al., 2024).
+[1], we define an MCQ $x$ is memorized by LLM $G$ if $G$ can extract the option contents $x_{O}$ exactly given the question $x_{Q}$.
+In practice, given the commonly used next-token prediction pre-training, we incorporate meta-information (e.g., benchmark name) and 5-shot examples to elicit the memorized following content  [^4] We aim to evoke the memorization of LLMs by prompting this information.
+For detailed reasons, please refer to Sec.
+IV-C., and use greedy decoding (i.e., temperature fixed to 0) during extraction [4, 3] (refer to Appendix A for the complete prompt).
+Using MMLU [6] as the evaluation benchmark, we divide the test set MCQs into memorized and non-memorized subsets, where the memorized subset consists of 909–982 questions (accounting for 6.5%–7.0% of the total 14,006) depending on the tested LLMs LlaMA2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B.
+The detailed statistics of questions across the memorized/non-memorized subsets are given in Tab.
+1 in the Appendix A.
+Figure 2: Model performance on memorized and non-memorized subsets of MMLU, where ‘0s’ and ‘5s’ stand for zero- and five-shot prompting, respectively.
 We then compute the accuracy (ACC) of tested LLMs by subsets as a proxy of model performance under different memorization conditions.
@@
 This finding challenges the commonly held assumption that memorization directly improves model performance.
-In addition, it also implies the coexistence of two distinct learning paradigms within LLMs, which we term rote memorization and genuine capability learning, respectively.
+In addition, it also implies the coexistence of two distinct learning paradigms or extremes within LLMs, which we term rote memorization and genuine capability learning, respectively.
+We also observe that the majority of memorized questions are those relatively simple, i.e., not in MMLU-PRO [5] (refer to Tab.
+1 in the Appendix A), which is a commonly recognized difficult subset of MMLU .
+Since MMLU-Pro’s curation prioritizes complex, model-challenging questions by excluding simplistic ones, the non-inclusion of our memorized questions in MMLU-Pro serves as objective evidence of their relative simplicity.
+This alignment with MMLU-Pro’s explicit filtering criterion validates our claim as reasonable, eliminating the possibility that the memorized quenstions are relatively difficult.
 However, our pre-investigation has its limitations.
-The binary classification of memorization potentially overlooks more nuanced forms of learning.
-Additionally, using ACC as the performance metric does not truly capture model capacity.
-We address these two issues in the following subsections, which then ensure a disentangle analysis of rote memorization and capability learning.
-3.2 Quantifying LLM Memorization
-For quantifying the memorization of LLMs, prior research (Shi et al., 2023; Zhang et al., 2024) suggests that outlier tokens, which exhibit higher generation probabilities, are more likely to be found in memorized samples.
-Building on this idea, we develop a metric that utilizes the bottom $K\%$ of token probabilities within the generated sequence as a measure of memorization.
-Formally, the memorization score $F_{m}(\overline{x},G)$ of LLM $G$ on text sequence $\overline{x}$ is computed as follows:
-$F_{m}(\overline{x},G)=\frac{1}{|\mathcal{M}_{K}(\overline{x})|}\sum\limits_{\overline{x}_{i}\in\mathcal{M}_{K}(\overline{x})}\log p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1}),$ (1)
-where $p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1})$ denotes the generation probability of token $\overline{x}_{i}$ by $G$ given its prefix subsequence as context, and set $\mathcal{M}_{K}(\overline{x})$ includes the $K\%$ of tokens with the lowest probabilities.
-The higher $F_{m}$ is, the more likely $\overline{x}$ is memorized by the LLM, i.e., the least memorized content could still been extracted with a high probability.
-3.3 Measuring LLM Capability with TrinEval
-We next present TrinEval, a novel evaluation framework designed to provide a more reliable measure of LLM performance by minimizing the influence of rote memorization.
-To understand how LLMs store and manipulate knowledge, Allen-Zhu and Li (2023) created a fictional biography dataset that enumerates various attributes (e.g.,, names, jobs, universities) and trained LLMs on this dataset.
-They employed a linear query-based probing method to uncover correlations between the entity token embeddings and the associated attributes, revealing that where LLMs encode knowledge, e.g., under person names or sequence of the knowledge mention, is crucial for robust mastery of knowledge.
-This insight leads us to believe that entity tokens, which should ideally store related knowledge, are the target for evaluating an LLM’s genuine capability.
+Using ACC as the performance metric does not truly capture model capacity.
+Additionally, the binary classification of memorization potentially overlooks more nuanced forms of learning due to the discrete classification cutoff.
+We address these two issues in the following Sec.
+III-B and Sec.
+III-C, respectively, which then ensure a disentangled and quantitative analysis of rote memorization and capability learning.
+III-B Measuring LLM Capability with TrinEval
+First, in order to truly capture model capacity, we present TrinEval, a novel evaluation framework designed to provide a reliable measure of LLM performance by minimizing the influence of rote memorization.
+TrinEval first preliminarily extract the knowledge-centric triplet of each question, detecting where the knowledge is stored and how to expose the knowledge through query-based probing.
+The framework then verify and reflect whether the trinity is qualified or not.
+Finally, TrinEval utilizes the trinity to evaluate the genuine capability of LLMs.
+To understand how LLMs store and manipulate knowledge, Allen et al.
+[18] created a fictional biography dataset that enumerates various attributes (e.g.,, names, jobs, universities) and trained LLMs on it.
+They employed a prompt-based linear probing method to uncover correlations between the entity token embeddings and the associated attributes, revealing that where LLMs encode/store knowledge, e.g., under person names or a sequence of knowledge augmentation, is crucial for robust mastery of knowledge.
+This insight leads us to believe that entity tokens, which should ideally store knowledge, are the target for evaluating an LLM’s genuine capability.
 However, applying this method to real-world datasets, such as MMLU, presents challenges.
@@
 TrinEval is a pluggable augmentation on any MCQ benchmarks and could expose the genuine capability of LLMs by verifying whether they have correctly encoded knowledge.
-We next explain the elements in the trinity and how to reformulate.
+In contrast to Allen et al.
+[18], TrinEval introduces a verbal query-based probing mechanism.
+It reformulates MCQs into a structured triplet format (entity–attribute–context), enabling direct assessment of knowledge encoded in entity tokens—without relying on model internals.
+This shift from embedding-space analysis to verbalized knowledge extraction is a conceptual and methodological advance.
+Generally, we instruct LLMs to extract the triplet according to the prompt and check and refine it if they are not qualified.
+This systematic validation is absent in most prior rephrasing methods and is empirically shown to be knowledge-preserving without introducing extraneous cues.
+The reformulation is completed by a two-round reflection-based prompting method, with the detailed procedure shown in Alg.
+1 and related prompts available in Fig.
+2 in Appendix B.
+Here, we explain the elements of the trinity and how to reformulate them.
 Knowledge entity.
-We suppose that if an LLM has mastered some knowledge, the key information pertinent to the knowledge should be encoded within a few subject tokens, namely knowledge entity, to support efficient retrieval.
+We suppose that if an LLM has mastered some knowledge, the key information pertinent to the knowledge should be encoded within a few subject tokens, namely, knowledge entity, to support efficient retrieval.
 By isolating these tokens, TrinEval ensures that only the essential information is considered.
 Attribute.
-The attribute acts as a verbal probe to guide the model focusing on the specific feature or property of the knowledge entity being inquired.
+The attribute acts as a verbal probe to guide the model to focus on the specific feature or property of the knowledge entity being inquired about.
 This mechanism allows TrinEval to isolate and assess the model’s understanding of the critical aspects of the questioning subject.
 Context.
-In a certain portion of questions, the conditions or background context can significantly influence the solution approach.
-By explicitly including context in the evaluation process, TrinEval helps the model account for relevant situational details that might otherwise be overlooked, ensuring that the model’s answer is based on a comprehensive understanding of the problem.
-By extracting the core and necessary question information in this trinity format, the reformulation by TrinEval is knowledge-preserving for the purposes of assessment.
-In the meanwhile, it completely destructs the original token sequence, effectively reducing the influence of memorization.
+For some questions, conditions or background contexts can significantly influence the solution approach.
+By explicitly including context in the evaluation process, TrinEval helps the model account for relevant situational details that might otherwise be overlooked, ensuring that the answer is based on a comprehensive understanding of the problem.
+Traditional rephrasing often relies on lexical variations (e.g., synonym replacement or sentence restructuring), which may still preserve memorization cues.
+In contrast, by extracting the core and necessary question information in this trinity format, the reformulation by TrinEval is knowledge-preserving for the purposes of assessment.
+In the meantime, it completely destroys the original token sequence, effectively reducing the influence of memorization.
 We will empirically verify these properties through experiments.
-The reformulation is completed by a two-round reflection-based prompting method, with detailed procedure (Alg.
-1) and related prompts available in Appendix B.
+Based on the framework, we locate which token the knowledge is stored within and expose the genuine capability of LLMs with the trinity.
 Given an MCQ $x=(x_{Q},x_{O},x_{W})$, it first queries a capable reformulation LLM to derive the knowledge entity $x_{E}$, attribute $x_{A}$, and context $x_{C}$ from the original $x$.
 The LLM is instructed that the triplet should be sufficient for answering the question correctly, without including the answer option itself, ensuring the integrity of the evaluation.
-The same LLM then assesses whether the triplet contains all necessary information and no redundant information (typically, the rote-memorization), in the meanwhile, yields a rationale $x_{L}$ as reflection (Shinn et al., 2024; Yao et al., 2022).
+In order to mimic the data contamination with in-context learning, we also present the original MCQ text in the format on the Huggingface dataset site as the “**Fact:**” part in the prompt, given that the potential data leakage is often caused by the data crawled on the Huggingface dataset site.
+The same LLM then assesses whether the triplet contains all necessary information and no redundant information (typically, the rote-memorization), meanwhile, yields a rationale $x_{L}$ as reflection [42, 41].
 If it does, the triplet is returned as the re-formulated question.
 Otherwise, the reformulation model refines the extraction, taking as input $x_{E}$, $x_{A}$, $x_{C}$, and $x_{L}$, and re-evaluates the updated triplet.
-Finally, prompting with the extracted $x_{E}$, $x_{A}$, and $x_{C}$ as well as options $x_{O}$, we inspect the generation probability of the ground-truth answer $x_{W}$ as the first token to measure capability:
-$F_{c}(x,G)=p_{G}(x_{W}|x_{E},x_{A},x_{C},x_{O}).$ (2)
-As can be seen, the $F_{c}$ metric retains the necessary knowledge-centric information while discarding unnecessary biases, especially the rote memorization of LLMs, which leads to the quantification of genuine capability of LLMs.
-4 Experiments
-In this section, we conduct extensive experiments to answer the following questions:
-Q1. Does TrinEval effectively preserve knowledge to enable accurate knowledge assessment?
-Q2. To what extent does TrinEval mitigate memorization effects during capability evaluation?
-Q3. What insights does TrinEval provide into the distinction between rote memorization and genuine capability in LLMs?
-4.1 Experiment Setup
-Models.
-We employ API-accessible commercial LLMs for question reformulation in TrinEval, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024).
-For model evaluation, we utilize open-source LLMs to obtain token-level logits.
-Our experiments focus on three widely adopted models: Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b).
-These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis.
-All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory.
-Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
-Benchmarks.
-We evaluate LLMs on the widely used MMLU benchmark (Hendrycks et al., 2020), in light of growing evidence of data contamination (Touvron et al., 2023) [^3] https://hitz-zentroa.github.io/lm-contamination/ in many recent LLMs on this data.
-MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities.
-We find duplicated MCQs across subjects on MMLU, and thus de-duplicate the dataset, resulting in a test set of 14,006 unique questions.
-Evaluation.
-With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions.
-For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metrics.
-4.2 Q1. Is TrinEval Knowledge-preserving?
-Figure 3: Knowledge-preserving validation of TrinEval reformulation.
-The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively.
-Best viewed in color.
-We begin by verify whether TrinEval’s reformulation preserves essential knowledge, a prerequisite for reliable knowledge assessment.
-Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance.
-Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding knowledge entities, attributes, and contexts using Qwen, and 4,645 MCQs with associated triplets using GPT.
-We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats.
-Example MCQs and the questioning prompt templates in both formats are provided in Table 2 in the Appendix; results are summarized in Fig.
-3.
-Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats, respectively.
-Notably, the majority of these are answered correctly in both the original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats.
-Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting that the TrinEval reformulation does not omit essential information.
-Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone.
-This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance.
-Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for accurate answering, satisfying the requirements for reliable capability evaluation.
-4.3 Q2. Can TrinEval Reduce Memorization?
-(a) Evoc. with Llama2 and dev-fsp.
-(b) Evoc. with Mistral and dev-fsp.
-(c) Evoc. with Vicuna and dev-fsp.
-(d) Evoc. with Llama2 and seq-fsp.
-(e) Evoc. with Mistral and seq-fsp.
-(f) Evoc. with Vicuna and seq-fsp.
-Figure 4: The results of memorization evocation (evoc.) under various dataset-related context, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively.
-In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question.
-These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general.
-However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
-In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization in LLMs, thereby isolating and revealing the models’ genuine capabilities under various circumstances.
-Following the methodology of Bordt et al.
-(2024), we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
-To evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset (see Appendix E).
-For this set of experiments, we focus on LLaMA, Mistral, and Vicuna, as their open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$.
-Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we use the $F_{m}$ value obtained from the original MCQ format without any additional context (i.e.,, the vanilla MCQ) as a reference baseline.
-We then visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
-Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction.
-Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al.
-(2022a), we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization.
-Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context.
-For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats.
-Results are illustrated in Fig.
-4.
-Consistent with the observations by Bordt et al.
-(2024), the results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects.
-However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context used.
-Notably, even under the strongest memorization-evoking setting, TrinEval’s absolute $F_{m}$ still stays below the vanilla baseline.
-These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling spurious recall from genuine model capability.
-4.4 Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability
-(a) $F_{m}$ v.s. $F_{c}$ with Llama on Qwen-extracted triplets.
-(b) $F_{m}$ v.s. $F_{c}$ with Mistral on Qwen-extracted triplets.
-(c) $F_{m}$ v.s. $F_{c}$ with Vicuna on Qwen-extracted triplets.
-(d) $F_{m}$ v.s. $F_{c}$ with Llama on GPT-extracted triplets.
-(e) $F_{m}$ v.s. $F_{c}$ with Mistral on GPT-extracted triplets.
-(f) $F_{m}$ v.s. $F_{c}$ with Vicuna on GPT-extracted triplets.
-Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs.
-capability metric $F_{c}$.
-According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs into 25 groups and visualize the MCQ distribution from weak to strong with heatmaps.
-In this subsection, we aim to explicitly examine the relationship between memorization and genuine problem-solving capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability).
-Since commercial API-based models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics.
-After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution.
-We then use heatmaps to visualize the distribution of samples across these groups, thereby revealing the relationship between memorization and capability for each tested model.
-As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap.
-This pattern is evident and consistent across our tests.
-For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value $<0.05$) for $F_{m}$ vs.
-$F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value $<0.05$).
-Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value $<0.05$), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value $<0.05$).
-We take the MCQs within the upper-right 2 X 2 squares in Fig.
-5 as the ones memorized by rote and compute the ratio of 19.6% by averaging the percentage of all evaluated LLMs.
-Similarly, 20.5% is computed as the ratio of MCQs within the lower-left 2 X 3 squares with genuine capability.
-Additional statistics results are reported in Table 3 in Appendix.
-These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall.
-To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) (Shiffrin, 2003).
-Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability.
-In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework.
-This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts.
-This dichotomy aligns with recent observations in LLMs.
-As shown by Allen-Zhu and Li (2023) and Ovadia et al.
-(2023), models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations.
-When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge.
-In this sense, LLMs are potential rote learners.
-Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation.
-Further experimental details and results are provided in Appendix C.
-Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings.
-‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$.
-Further results are shown in Appendix F.
-To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% of samples.
-Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus.
-As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs.
-By selecting only the top 1% nearest neighbors, we aim to retain the samples that are semantically-correlated in knowledge description.
-We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions of the heatmap in Fig.
-5.
-The corresponding results are presented in Fig.
-6.
-Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs.
-This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters.
-This observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations.
-The results of alternative thresholds are shown in Table 5.
-Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.
-5 Conclusion
-This study provided a novel perspective on benchmark contamination in LLM evaluation, reframing it as an inherent aspect of learning.
-This perspective led us to explore the relationship between memorization and genuine capability in LLMs.
-Through our empirical investigation, we observed a surprising result: LLMs performed worse on memorized MCQs compared to those not, suggesting that superficial memorization may undermine problem-solving ability rather than enhance it.
-This finding also implies the existence of two distinct learning paradigms in LLMs: rote memorization and genuine capability learning.
-To disentangle them, we proposed TrinEval, a novel evaluation method that reformulates MCQs into a knowledge-centric trinity, thus separating the influence of memorization from genuine knowledge application.
-Experiments validated both the knowledge-preserving and memorization-reducing properties of this approach.
-Based on that, TrinEval reveals the in-robustness of LLMs’ knowledge learning, e.g., popular open-source LLMs memorize 20.5% of knowledge points by rote without understanding in MMLU.
-As such, we believe this work lays the groundwork for future studies on improving LLM knowledge robustness and more thorough evaluation.
-6 Limitations
-Our limitations are mainly two points.
-First, though our proposed TrinEval retrains the problem-solving ability of the LLMs and obtains stronger robustness, it is not a dynamical re-organizing method that can still be leaked and pre-experienced during training.
-On the one hand, we appeal to the LLM developers not to use this re-organizing method as part of the training corpus.
-On the other hand, future works will be focused on developing dynamic evaluation method (Zhu et al., 2023, 2024).
-Second, we did not give a clear exploration on how and why the more LLMs memorize, the less the capability of the LLMs obtains.
-In future work, we will also look into the mechanism of the training and structure of LLMs for a thorough study of the phenomenon.
-Appendix A Details of the extracting prompts and the extracted (memorized) MCQs
-In this section, we introduce the details of the processed dataset and the prompts for extraction.
-MCQs from some subjects contain similar or identical options [^4] E.g., the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’)..
-With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio.
-In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset.
-After deduplication, we obtain 14,006 MCQs for evaluation.
-The extraction prompt and the detailed statistics are shown in the following text and Tab.
-1.
-Prompt template for extraction: Prompt template for pre-investigation on LLM Memorization w.r.t. Capability:
-You are an expert of multiple choice questions of MMLU dataset.
-The following are multiple-choice questions (with answers) about [subject].
-[examples] [question] Options: A.
-You are an expert of multiple choice questions of MMLU dataset.
-The following are multiple choice questions (with answers) about [subject].
-[examples] [question] Options: A.
-[content for option A] B.
-[content for option B] C.
-[content for option C] D.
-[content for option D] Answer:
-{tblr}
-cells = c, cell21 = r=3, cell51 = r=3, cell81 = r=3, vline2-3 = 1-2,5,8, vline2-3 = 3-4,6-7,9-10, hline1-2,5,8,11 = -, Model & Subset Simple Pro MMLU
-Llama memorized 912 70 982
-non-mem. 6,548 6,476 13,024
-all 7,460 6,546 14,006
-Mistral memorized 879 36 915
-non-mem. 6,581 6,510 13,091
-all 7,460 6,546 14,006
-Vicuna memorized 893 16 909
-non-mem. 6,567 6,530 13,097
-all 7,460 6,546 14,006
-Table 1: Statistics of memorized and non-memorized questions by Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B in MMLU.
-Appendix B Details of TrinEval
-In this section, we introduce the details of the proposed TrinEval.
-The pseudo-code is shown in the Alg.
-1.
-The prompts used are also shown below.
-Note that the potential data leakage is often caused by the data crawled on the Huggingface dataset site.
-Thus, we also provide the original text of MCQs in the format on the Huggingface dataset site to mimic the data contamination with in-context learning.
+The details of the reorganization algorithm are shown as follows.
+Reflection prompts are shown in Appendix B.
 Algorithm 1 MCQ reformulation by TrinEval
@@
 12:  else
-13:        Discard the MCQ, return $None$;
+13:   Discard the MCQ, return NoneNone;
 14:  end if
 15:end if
-Prompt template for triplet extraction:
-You are an expert of Knowledge Keyword extraction.
+Finally, prompting with the extracted $x_{E}$, $x_{A}$, $x_{C}$, and options $x_{O}$ (see Fig.
+5 in Appendix B for illustration), we inspect the generation probability of the next ground-truth answer token $x_{W}$ (i.e., A/B/C/D) as the measurement:
+$F_{c}(x,G)=p_{G}(x_{W}|x_{E},x_{A},x_{C},x_{O}).$ (1)
+As can be seen, $F_{c}$ metric retains the necessary knowledge-centric information while discarding unnecessary biases, especially the rote memorization of LLMs, which leads to the quantification of genuine capability of LLMs.
+III-C Quantifying LLM Memorization
+On the other hand, in order to look into the nuanced forms of learning, we believe that a continuous metric is required for memorization evaluation.
+For quantifying the memorization of LLMs, prior research [16, 17] suggests that outlier tokens, which exhibit higher generation probabilities [^5] As established in Shi et al.
+[16]: “an unseen example is likely to contain a few outlier words with low probabilities under the LLM, while a seen example is less likely to have words with such low probabilities.”, are more likely to be found in memorized samples.
+Experiments on WIKIMIA [16] also finds that this method exhibits high AUC score on seen samples.
+In order to avoid discrete binary classification, building on this idea, we adopt the metric that utilizes the bottom $K\%$ of token probabilities within the generated sequence as a measure of memorization (we primarily set the value of $K$ to 10 in this paper).
+Formally, the memorization score $F_{m}(\overline{x},G)$ of LLM $G$ on text sequence $\overline{x}$ is computed as follows [^6] In the experiment, we have also tested other similar metrics like Min-K++% [17], which show consistent results (exhibit a Pearson correlation coefficient of 0.9007 (strong positive correlation)).
+Thus, we only use Min-K% [16] as an example here:
+$F_{m}(\overline{x},G)=\frac{1}{|\mathcal{M}_{K}(\overline{x})|}\sum\limits_{\overline{x}_{i}\in\mathcal{M}_{K}(\overline{x})}\log p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1}),$ (2)
+where $p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1})$ denotes the generation probability of token $\overline{x}_{i}$ by $G$ given its prefix subsequence as context, and set $\mathcal{M}_{K}(\overline{x})$ includes the $K\%$ of tokens with the lowest probabilities.
+The higher $F_{m}$ is, the more likely $\overline{x}$ is memorized by the LLM, i.e., the least memorized content could still been extracted with a high probability.
+Note that our objective is not to achieve high-precision quantification of memorization.
+Instead, we aim to statistically demonstrate a valid quantification of memorization across vast samples for comparative analysis.
+IV Experiments
+In this section, we conduct extensive experiments to answer the following questions:
+Q1. Does TrinEval effectively preserve knowledge to enable accurate knowledge assessment?
+Q2. To what extent does TrinEval mitigate memorization effects during capability evaluation?
+Q3. What insights does TrinEval provide into the rote memorization and genuine capability learning in LLMs?
+IV-A Experiment Setup
+Models.
+We employ proprietary commercial LLMs for question reformulation in TrinEval, i.e., gpt-4o-2024-08-06(GPT) [13], qwen-max-2024-09-19(Qwen) [15, 14], deepseek-r1(Dpsk) [50] and qwen3-max-2026-01-23(Qwen3) [51].
+For model evaluation, we utilize open-source LLMs to obtain token-level logits.
+Our experiments first focus on three widely adopted models: LlaMA2-7B (LlaMA) [7], Mistral-7B-v0.2 (Mistral) [11], and Vicuna-v1.5-7B (Vicuna) [12] due to their reported data contamination evidence [3, 7] on MMLU.
+Then, we further extend the results to the latest larger models, such as LlaMA3.1-8B [52], Qwen3-14B, and Qwen3-32B.
+These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis.
+All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory.
+Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
+Benchmarks.
+We evaluate LLMs on the widely used MMLU benchmark [6], in light of growing evidence of data contamination [7] in many recent LLMs on this data.
+MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities.
+We de-duplicate MCQs across subjects on MMLU.
+Also, MCQs from some subjects contain similar or identical options777e.g., the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’)..
+With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio.
+In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset, resulting in a test set of 14,006 unique questions.
+We also extend the evaluation on the widely used GSM8K benchmark [53], given the determinacy of its answers and the verified data contamination [3].
+GSM8K consists of high-quality grade school math problems, focusing on multi-step arithmetic reasoning to rigorously evaluate the basic mathematical problem-solving ability of LLMs.
+In this way, we are able to evaluate the memorization of the OEQs according to the extraction of the official chain-of-thought(COT) sequences.
+We mainly select the test set of GSM8K as the evaluation benchmark.
+After filtering the duplicate OEQs, we obtain an OEQ set of 1,319 unique questions.
+Evaluation.
+With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions on MCQs.
+For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metric for MCQs.
+For OEQs, we compute the perplexity of the answer string as the capability metric, only available with the open-source LLMs.
+IV-B Q1. Is TrinEval Knowledge-preserving?
+We begin by verify whether reformulation of TrinEval preserves essential knowledge, a prerequisite for reliable knowledge assessment.
+Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance.
+We verified the above two objectives through both automatic analysis based on LLMs’ results and manual annotations, ensuring the robustness of TrinEval.
+Q1.1 Automatic analysis
+Figure 3: Knowledge-preserving validation of TrinEval.
+The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively.
+Best viewed in color.
+Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding triplets using Qwen, and 4,645 MCQs with associated triplets using GPT.
+We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats.
+Example MCQs and the questioning prompt templates in both formats are provided in Fig.
+5 in Appendix B; results are summarized in Fig.
+3.
+Here, we primarily employ robust proprietary LLMs to evaluate the knowledge retention capability of the TrinEval framework.
+This approach effectively mitigates potential interference from the models’ memorization effects, thereby ensuring accurate assessment of their intrinsic knowledge processing abilities and preventing erroneous conclusions that might arise from conflating memorization with genuine comprehension.
+Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats.
+Notably, the majority of these are correctly answered in both original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats.
+Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting the TrinEval reformulation does not omit essential information.
+Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone.
+This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance.
+Q1.2 Manual annotation
+Given the potential risk for the LLM-based Knowledge Preserving evaluation (line 4 and line 10 in 1), the proprietary LLMs might still be able to answer the MCQs without sufficient knowledge.
+Since this is still prompting LLMs who have been trained on these datasets and “know” the original content, human annotation is also applied.
+The annotation of each MCQ encompasses three subtasks: (1) answering the question in the re-organized form, (2) answering the question in the original form, and (3) verifying if the reformulation is Knowledge Preserving or not.
+Metric Qwen GPT
+TrinEval correct only 0.0% 0.0%
+both correct 96.296% 96.667%
+original correct only 3.704% 3.333%
+K.P. score 4.101 4.369
+TABLE I: Result of Human Annotation on LLM-based knowledge preserving (K.P.) evaluation in TrinEval
+For efficient annotation, we implemented a stratified sampling procedure by selecting one MCQ per subject from all 56 MMLU subjects (as a temporary compromise for limited time, which will be expanded later) under both Qwen and GPT reorganization paradigms.
+This yielded 112 representative questions (2 systems × 56 subjects) for evaluation.
+Three human annotators independently performed dual-form assessments through: (1) Direct question answering with the reformed format first and the original format; (2) Knowledge Preservation (K.P.) scoring across two dimensions: i.
+Knowledge adequacy (sufficiency for accurate response), ii.
+removal of redundant content using a 5-point scale (1=unsatisfactory; 2=major information is missed or unnecessary information is incorporated, but part is still acceptable, 3=need to take some time to understand, but can still solve the MCQ, 4=an element properly belonging to one triplet component appears in another, but does not impact MCQ solving; 5=optimal).
+We use a continuous rather than a binary metric to mitigate the cognitive difference in the threshold between the annotators.
+Inter-rater reliability was ensured through consensus-building discussions prior to formal annotation.
+Final scores were aggregated using averaged values to further mitigate individual annotator bias.
+The results are shown in Tab.
+I below.
+Our analysis reveals that over 95% of correctly answered MCQs maintained consistency across both original and paraphrased formats.
+Furthermore, human annotators rated our paraphrased questions mean K.P.
+score exceeding 4.0 (on a 5-point scale), which means that the reformulated MCQs only somehow influence the readability of humans but do not impact the solvability of the original format.
+This provides empirical validation that our proposed TrinEval methodology effectively preserves necessary knowledge elements from original MCQ formulations, while the influence of the LLM memorization during the evaluation is rather limited.
+Experimental results under human annotation also reveal that Qwen underperforms GPT in key metrics, particularly in processing long-context texts, where it occasionally omits background information (evidenced by excessive ”N/A” assignments in the Context fields).
+This capability gap is further reflected in MCQ annotations: there are only MCQs that are merely correctly answered with the original format, except for the correct MCQs with both formats.
+Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for answering, satisfying the requirements for reliable capability evaluation.
+IV-C Q2. Can TrinEval Reduce Memorization?
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Figure 4: Results of memorization evocation (evoc.) under dataset-related contexts, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively.
+In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question.
+These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general.
+However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
+In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization, thereby isolating and revealing the LLMs’ genuine capabilities under various circumstances.
+Following Bordt et al.
+[4], we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
+We suppose that unintentional data contamination arises from crawling dataset pages (e.g., Hugging Face) during the compilation of LLM pretraining datasets.
+When researchers are organizing the pretraining corpus, one or more neighboring original data samples would be truncated and concatenated sequentially into a pretraining sample.
+Thus, according to Carlini’s theory [1] and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
+Thus, following the previous study [4], to evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset, to find out if the reformulation of the TrinEval framework can eliminate the memorization effect.
+For this set of experiments, we focus on LlaMA, Mistral, and Vicuna, as there exists verified evidence of data contamination and open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$.
+Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we take the $F_{m}$ value of the original MCQ format without any additional context (i.e.,, vanilla MCQ) as a reference baseline.
+We visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
+Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction.
+Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al.
+[1], we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization.
+Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context.
+For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats.
+Results are illustrated in Fig.
+4.
+Consistent with the observations by Bordt et al.
+[4], results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects.
+However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context.
+Notably, even under the strongest memorization-evoking setting, TrinEval ’s absolute $F_{m}$ still stays below the vanilla baseline.
+These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling recall from genuine model capability.
+IV-D Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability Learning
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs.
+capability metric $F_{c}$.
+According to the values of $F_{m}$ and $F_{c}$, we separate MCQs into 25 groups and visualize the distribution from weak to strong with heatmaps.
+Same as below.
+In this subsection, we aim to examine the relationship between memorization and genuine capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability).
+Q3.1 TrinEval’s findings on MCQs
+Since commercial proprietary models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics.
+After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution.
+We then use heatmaps to visualize the distribution of samples across these groups, revealing the relationship between memorization and capability for each tested model.
+As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap.
+For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value <0.05<0.05) for $F_{m}$ vs.
+$F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value <0.05<0.05).
+Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value <0.05<0.05), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value <0.05<0.05).
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Figure 6: The distribution of MCQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$ with latest larger models.
+We analyze the distribution in detail by revealing the ratio of MCQs within the upper right and lower left 2 $\times$ 2 and 3 $\times$ 3 squares as well as the Pearson correlations between the $F_{m}$ and $F_{c}$ of these MCQs.
+We take the MCQs within the lower-left 2 $\times$ 2 squares in Fig.
+5 as the ones memorized by rote and compute the ratio of 20.5% by averaging the percentage of all evaluated LLMs.
+Similarly, 19.6% is computed as the ratio of MCQs within the upper-right 2 $\times$ 2 squares with genuine capability.
+All of these MCQs also exhibit a tendency towards a negative correlation between the capabilities and memorization of LLMs (all Pearson correlation values are computed with p-values at the 0.05 level).
+Additional statistics results are reported in Tab.
+III in Appendix C.
+Further, we also plot the curve of the separation of the rote memorization and genuine capability learning MCQs in Appendix C.
+With each progressing $F_{m}$ as threshhold, the tested MCQs exhibit prominent $F_{c}$ gaps between the two groups, showing significant difference in learning extremes.
+These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall.
+Still, we would argue that knowledge acquisition in LLMs is likely a continuous spectrum rather than a strict binary.
+Our preliminary investigation also highlighted the limitations of a discrete classification, and thus, we apply the continuous quantification metric $F_{m}$ and $F_{c}$.
+As seen in the heatmaps, some questions fall in intermediate regions, indicating varying degrees of memorization and understanding.
+Thus, while we focus on the two extremes for analytical clarity, we agree that other intermediate or hybrid categories may exist.
+Q3.2 Can TrinEval’s findings be extended to latest larger LLMs?
+Based on previous research [7, 3], the aforementioned experiments were conducted in a relatively deterministic experimental environment, i.e., confirmed data contamination evidence.
+Therefore, we attempt to extend the above results to the latest larger LLMs.
+Here, we mainly reformulate the MMLU MCQs based on Dpsk and Qwen3, obtaining 6,115 and 5,905 MCQs, respectively, and verify the results based on LlaMA3.1-8B, Qwen3-14B, and Qwen3-32B.
+The experimental procedure remains identical to that described in the preceding experiment.
+Results are shown in Fig.
+6.
+We can see that the majority of MCQs also cluster in the lower-left and upper-right corners of the heatmap, which is evident and consistent across our tests.
+For instance, in the results of LlaMA3.1 using Qwen3-extracted triplets, 38.34% of the MCQs fall within the two opposite corner regions of the $2\times 2$ grid.
+For the examples in these two regions, the Pearson correlation between $F_{m}$ and $F_{c}$ is -0.7771 (p-value <0.05<0.05).
+Extending the analysis to the $3\times 3$ corner regions increases coverage to 78.34%, yielding a correlation of 0.5506 (p-value <0.05<0.05).
+Similarly, for Qwen-14B on Qwen triplets, 39.04% of the MCQs lie in the two $2\times 2$ corner regions, with a correlation of -0.7260 (p-value <0.05<0.05); and 79.02% fall in the two $3\times 3$ corner regions, with a correlation of -0.4905 (p-value <0.05<0.05).
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Figure 7: The distribution of OEQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$.
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Figure 8: The distribution of OEQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$ with latest larger models.
+Q3.3 Can TrinEval’s findings be extended to OEQs?
+Subsequently, we pose a new question: can the above findings on MCQs be extended to OEQs?
+To answer this question, we apply TrinEval on the GSM8K dataset.
+For $F_{m}$, instead of the original option content in MCQs, we select the bottom K% token probabilities of the official CoT seuqence for each OEQ to evluate the memorization of LLMs.
+As to $F_{c}$, given that GSM8K is a mathematics problem benchmark with a definite answer, we utilize the perplexity of the answer sequence.
+For open-ended questions (OEQs) that admit multiple correct answer formulations and require semantic correctness evaluation, we believe that our TrinEval can also be readily adapted to corresponding measurement.
+We utilize TrinEval to reformulate the OEQs and obtain 1282 qualified questions and use them to evaluate LlaMA, Mistral, Vicuna, as well as latest larger models, Llama3.1-8B, Qwen3-14B, and Qwen3-32B.
+The results are shown in Fig.
+7 and Fig.
+8.
+We surprisingly find that on OEQs, most LLMs still exhibit the same phenomena, i.e., OEQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance.
+Among them, for LLaMA on the triplet questions extracted by Dpsk, 18.04% of the questions exhibited rote memorization, while 17.62% were genuinely mastered.
+For Qwen3‑32B on the same Dpsk‑extracted triplets, 17.62% of the questions showed rote memorization, and 17.53% were genuinely mastered.
+However, for some results, for instance, LLaMA3.1‑8B on the triplet extracted by Qwen3, only 15.76% of the questions were classified as rote memorization, falling below the 16% baseline in the $5\times 5$ heatmap.
+We attribute this to the following reasons.
+First, comparing Fig.
+7 and Fig.
+8, we observe that the distinction between rote memorization and genuine capability learning is less pronounced in the latest larger LLMs than in other LLMs.
+This may be because, as suggested by recent research [55, 56, 21], LLMs are increasingly designed to be trained on reformulated corpus during pretraining, thereby reducing the formation of such rote memorization.
+Second, Several studies [54] have also found that overly formatted corpora can lead to poorer generalization in LLMs.
+Thus, compared to MCQs, OEQs are less structured and thus may be less susceptible to rote memorization.
+Nevertheless, the conclusions drawn from most LLMs remain consistent with our earlier experimental findings on MCQs, suggesting that LLMs also exhibit rote memorization in the context of OEQs.
+Q3.4 How to interpret TrinEval’s findings?
+To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) [46].
+Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability.
+In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework.
+This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts.
+This dichotomy aligns with recent observations in LLMs.
+As shown by Allen et al.
+[18] and Ovadia et al.
+[21], models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations.
+When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge.
+In this sense, LLMs are potential rote learners.
+Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation.
+Figure 9: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings.
+‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits a high memorization metric $F_{m}$ and a low capability $F_{c}$ while ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$.
+To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% samples.
+Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus.
+As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs.
+As there are 57 different subjects within the MMLU dataset, we argue that unrelated sequences would lead to increased embedding distance even though they are among the mastered knowledge points.
+Here, we try to filter out the unrelated samples for each MCQ and thus keep the closest samples at the 1e-2 level in order to make sure there are not too many unrelated samples incorporated.
+To highlight this result more prominently, we employ a relatively stringent filtering strategy and made it 1% of the closest samples, aiming to retain the samples that are semantically correlated in knowledge description.
+We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions in Fig.
+5.
+The corresponding results are presented in Fig.
+9, where RM stands for rote memorization, and GCL stands for genuine capability learning.
+In order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5% in Tab.
+II in Appendix D.
+Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs.
+This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters.
+The observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations.
+Also, we can see that the more samples we incorporated, the higher the average distance of the closest embeddings grows.
+Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings.
+As a conclusion, though it is widely believed that memorization may lead to better but cheating performance of LLMs, we find that the more LLMs memorize, the worse they are at solving problems.
+V Conclusion
+This study provided a novel perspective on benchmark contamination in LLM evaluation, reframing it as an inherent aspect of learning.
+This perspective led us to explore the relationship between memorization and genuine capability in LLMs.
+Through our empirical investigation, we observed a surprising result: LLMs performed worse on memorized benchmarks compared to those not, suggesting that superficial memorization may undermine problem-solving ability rather than enhance it.
+This finding also implies the existence of two distinct learning paradigms in LLMs: rote memorization and genuine capability learning.
+To disentangle them, we proposed TrinEval, an evaluation method reformulating MCQs and OEQs into a knowledge-centric trinity and separating the influence of memorization from genuine knowledge application.
+Experiments validated both knowledge-preserving and memorization-reducing properties of TrinEval.
+Based on that, TrinEval reveals the in-robustness of LLMs’ knowledge learning, e.g., popular open-source LLMs memorize 20.5% of knowledge points by rote in MMLU.
+We also discussed the generalizability of our approach to other forms of questions (e.g., open-ended questions) in Sec IV-C.
+As such, we believe this work lays the groundwork for future studies on improving LLM knowledge robustness and a more thorough evaluation.
+Appendix A Details of the extracting prompts and the extracted (memorized) MCQs
+In this section, we elaborate on the details of the processed dataset and the prompts for extraction.
+MCQs from some subjects contain similar or identical options888e.g.,, the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’)..
+With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio.
+In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset.
+After deduplication, we obtain 14,006 MCQs for evaluation.
+The detailed prompt and statistics are shown in the following Fig.
+10 and Tab.
+II.
+Model Subset Simple Pro MMLU
+Llama memorized 912 70 982
+non-mem. 6,548 6,476 13,024
+all 7,460 6,546 14,006
+Mistral memorized 879 36 915
+non-mem. 6,581 6,510 13,091
+all 7,460 6,546 14,006
+Vicuna memorized 893 16 909
+non-mem. 6,567 6,530 13,097
+all 7,460 6,546 14,006
+TABLE II: Statistics of memorized and non-memorized questions by Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B in MMLU.
+Appendix B Details of TrinEval
+In this section, we introduce the details of the reflection part of the proposed TrinEval.
+The extraction and reflection prompts used are shown in Fig.
+11, Fig.
+12, and Fig.
+13 below.
+The reformulated example is shown in Fig.
+14.
+Appendix C Detailed results of memorization v.s. capability
+In this section, we exhibit the detailed results of the Q3. What does TrinEval reveal about the memorization v.s. the capability of LLMs, as shown in Tab. IV.
+Further, inspired by the Precision-Recall Curve, we take each unique $F_{m}$ of the qualified MCQs as the threshold to separate them as the Memorized and Capable MCQs.
+For each separation, we compute the probability of whether the $F_{c}$ of a randomly selected Capable MCQ exceeds the $F_{c}$ of a randomly selected Memorized MCQ and plot them as the blue curve.
+We also compute the T-test p-value between the $F_{c}$s of the Memorized MCQs and Capable MCQs as the green curve.
+The results are shown in Fig.
+15.
+For the second row, we filter out the MCQs within the upper left and lower right $2\times 2$ squares.
+From the figure, we observe that over a relatively long segment in the middle of the x-axis threshold range, the probability remains at a comparatively high value, while the p-value stays below 0.05.
+From this, we can conclude that $F_{m}$ can distinguish between MCQs with high $F_{c}$ and those with low $F_{c}$ with a negative correlation at a high confidence level.
+Appendix D Detailed results of Q3.4
+To highlight this result more prominently, we employed a relatively stringent data filtering strategy in the paper and made it 1% of the closest samples in Sec.
+IV-D.
+In this section, in order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5%.
+The results are shown in Tab.
+III, where RM stands for rote memorization, and GCL stands for genuine capability learning.
+Threshold 1% 3% 5%
+Subset RM GCL RM GCL RM GCL
+Llama-Qwen 29.538 16.501 32.718 17.281 33.790 18.060
+Llama-GPT 28.925 17.777 31.327 18.275 32.315 19.068
+Mistral-Qwen 78.663 47.648 86.003 52.995 89.598 55.525
+Mistral-GPT 82.932 56.375 90.597 62.112 94.384 64.820
+Vicuna-Qwen 38.761 21.006 41.396 23.189 42.490 24.185
+Vicuna-GPT 37.037 22.996 41.396 23.189 42.490 24.185
+TABLE III: Result of averaged embedding distance of the closest MCQs under different thresholds
+We can see that, as we concluded in the main part, the more samples we incorporated, the higher the average distance of the closest embeddings grows.
+Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings.
+This proves that the reported results are robust and solid.
+Prompt template for extraction:
+You are an expert of multiple choice questions of the MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A.
+Prompt template for pre-investigation on LLM Memorization w.r.t.
+Capability: You are an expert of multiple choice questions of the MMLU dataset.
+The following are multiple-choice questions (with answers) about [subject].
+[examples] [question] Options: A.
+[content for option A] B.
+[content for option B] C.
+[content for option C] D.
+[content for option D] Answer:
+Figure 10: Instructions for memorization extraction and model performance evaluation, respectively.
+LLMs Dataset $2\times 2$ squares $3\times 3$ squres
+Ratio (%) Pearson correlation Ratio (%) Pearson correlation
+LlaMA2-Qwen All 38.57 -0.7755 74.17 -0.8124
+Simple 37.07 -0.7784 72.63 -0.8121
+Pro 38.66 -0.783 74.51 -0.8109
+LlaMA2-GPT All 35.22 -0.7835 71.04 -0.7924
+Simple 33.9 -0.777 69.62 -0.7919
+Pro 35.45 -0.7916 71.54 -0.7881
+Mistral-Qwen All 44.9 -0.8722 80.82 -0.8794
+Simple 38.47 -0.8494 74.04 -0.8271
+Pro 44.32 -0.8045 80.08 -0.8682
+Mistral-GPT All 40.37 -0.8042 76.58 -0.8736
+Simple 35.51 -0.8297 72.27 -0.8664
+Pro 38.52 -0.7103 74.91 -0.7969
+Vicuna-Qwen All 42.94 -0.8771 79.23 -0.8365
+Simple 37.86 -0.758 73.85 -0.7168
+Pro 42.01 -0.8609 77.86 -0.886
+Vicuna-GPT All 38.69 -0.8621 74.83 -0.8672
+Simple 34.77 -0.8096 70.71 -0.7775
+Pro 37.37 -0.7794 73.98 -0.8728
+TABLE IV: The ratio and the Pearson-correlation between the $F_{c}$ and $F_{m}$ of the MCQs within the upper right and lower left $2\times 2$ and $3\times 3$ squares.
+For LLMs, ‘LlaMA2-Qwen’ refers that the $F_{c}$ and $F_{m}$ are calculated with LlaMA2 based on the Qwen-extracted triplet, and similarly hereinafter.
+For the Dataset column, ‘All’ stands for all the qualified MCQs after the triplet extraction, ‘Pro’ refers to the qualified MCQs that are the members of the mmlupro dataset while ‘Simple’ refers to the rest of the MCQs that are relatively easier.
+Prompt template for triplet extraction You are an expert of Knowledge Keyword extraction.
 Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question.
@@
 **Fact:** [question]  [option content list]  [subject]  [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
-Prompt template for triplet validation & reflection:
-You are an expert of [subject] and an advanced reasoning agent that can determine whether the given Knowledge Keyword, Attribute of the Knowledge Keyword and the Context present most of the necessary information of the Question for obtaining the given Answer.
+Figure 11: Instruction for LLMs to extract the triplet.
+Prompt template for triplet validation & reflection You are an expert of [subject] and an advanced reasoning agent that can determine whether the given Knowledge Keyword, Attribute of the Knowledge Keyword and the Context present most of the necessary information of the Question for obtaining the given Answer.
 Suppose you have sufficient background knowledge about subj.
@@
 - **Question:** [question] - **Answer:** [answer] - **Knowledge Keyword:** [extracted knowledge entity] - **Attribute:** [extracted attribute] - **Context:** [extracted context]
-Prompt template for the second round triplet extraction:
-You are an advanced reasoning agent that can improve through self-reflection and an expert of Knowledge Keyword extraction.
+Figure 12: Instruction for LLMs to verify whether the triplet is qualified and reflect.
+Prompt template for the second round triplet extraction You are an advanced reasoning agent that can improve through self-reflection and an expert of Knowledge Keyword extraction.
 Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question.
@@
 **Fact:** [question]  [option content list]  [subject]  [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
-Original MCQ Original MCQ Example
-You are an expert on multiple choice questions of [subject].
+Figure 13: Instruction for LLMs on the second round triplet extraction.
+Trineval Reformulation Original MCQ You are an expert on multiple choice questions of [subject].
 Analyze the given question and the given options.
@@
 [option C] D.
-[option D] You are an expert on multiple choice questions of high school computer science.
+[option D] Original MCQ Example You are an expert on multiple choice questions of high school computer science.
 Analyze the given question and the given options.
@@
 The return address D.
-Stack pointer for the calling activation record
-TrinEval MCQ TrinEval MCQ Example
-You are an expert on multiple choice questions of [subject].
+Stack pointer for the calling activation record TrinEval MCQ You are an expert on multiple choice questions of [subject].
 Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question.
@@
 [option C] D.
-[option D] You are an expert on multiple choice questions of high school computer science.
+[option D] TrinEval MCQ Example You are an expert on multiple choice questions of high school computer science.
 Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question.
@@
 Stack pointer for the calling activation record
-Table 2: Template and an example of the Original MCQ template and the TrinEval MCQ template. [$\cdot$] refers to the blank that should be filled according to the content of each MCQ.
-Appendix C Detailed results of memorization v.s. capability
-In this section, we exhibit the detailed results of the Q3.
-What does TrinEval reveal about the memorization v.s.
-the capability of LLMs.
-We reveal the ratio of MCQs within the upper right and lower left 2 $\times$ 2 and 3 $\times$ 3 squares as well as the Pearson correlations between the $F_{m}$ and $F_{c}$ of these MCQs.
-Our analysis reveals a tendency towards a negative correlation between the capabilities and memorization of LLMs shown in the Tab.
-3.
-All person correlation values are computed with p-values at the 0.05 level.
-(a) Probability and p-value with Llama2 based on GPT.
-(b) Probability and p-value with Mistral based on GPT.
-(c) Probability and p-value with Vicuna based on GPT.
-(d) Probability and p-value with Llama2 based on GPT (21.59 % MCQs filtered).
-(e) Probability and p-value with Mistral based on GPT (15.95 % MCQs filtered).
-(f) Probability and p-value with Vicuna based on GPT (17.59 % MCQs filtered).
+Figure 14: Template and an example of the Original MCQ template and the TrinEval MCQ template. $\cdot$] refers to the blank that should be filled according to the content of each MCQ.
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
 Figure 15: The over-performing probability curve and p-value curve with different $F_{m}$ thresholds.
-In this figure, we take each unique $F_{m}$ as the threshold to separate the qualified MCQs as the Memorized and Capable MCQs.
-We compute the probability of a randomly selected Capable MCQ’s $F_{c}$ exceeds a randomly selected Memorized MCQ’s $F_{c}$ under each threshold as the blue curve, and the green curve is the p-value of the T-test between the $F_{c}$s of the Capable MCQs and the Memorized MCQs.
-Further, inspired by the Precision-Recall Curve, we take each unique $F_{m}$ of the qualified MCQs as the threshold to separate them as the Memorized and Capable MCQs.
-For each separation, we compute the probability of whether the $F_{c}$ of a randomly selected Capable MCQ exceeds the $F_{c}$ of a randomly selected Memorized MCQ and plot them as the blue curve.
-We also compute the T-test p-value between the $F_{c}$s of the Memorized MCQs and Capable MCQs as the green curve.
-The results are shown in Fig.
-7.
-For the second row, we filter out the MCQs within the upper left and lower right $2\times 2$ squares.
-From the figure, we observe that over a relatively long segment in the middle of the x-axis threshold range, the probability remains at a comparatively high value, while the p-value stays below 0.05.
-From this, we can conclude that $F_{m}$ can distinguish between MCQs with high $F_{c}$ and those with low $F_{c}$ with a negative correlation at a high confidence level.
-This further supports that LLMs are potential rote learners, the more the LLMs memorize, the more poorly they perform.
-{tblr}
-row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=2c, cell15 = c=2c, cell31 = r=3, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell61 = r=3, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell91 = r=3, cell93 = c, cell94 = c, cell95 = c, cell96 = c, cell103 = c, cell104 = c, cell105 = c, cell106 = c, cell113 = c, cell114 = c, cell115 = c, cell116 = c, cell121 = r=3, cell123 = c, cell124 = c, cell125 = c, cell126 = c, cell133 = c, cell134 = c, cell135 = c, cell136 = c, cell143 = c, cell144 = c, cell145 = c, cell146 = c, cell151 = r=3, cell153 = c, cell154 = c, cell155 = c, cell156 = c, cell163 = c, cell164 = c, cell165 = c, cell166 = c, cell173 = c, cell174 = c, cell175 = c, cell176 = c, cell181 = r=3, cell183 = c, cell184 = c, cell185 = c, cell186 = c, cell193 = c, cell194 = c, cell195 = c, cell196 = c, cell203 = c, cell204 = c, cell205 = c, cell206 = c, hline1,3,21 = -, hline2 = 3-6, LLMs & Dataset $2\times 2$ squares $3\times 3$ squres
-Ratio (%) Pearson correlation Ratio (%) Pearson correlation
-Llama2-Qwen All 38.57 -0.7755 74.17 -0.8124
-Simple 37.07 -0.7784 72.63 -0.8121
-Pro 38.66 -0.783 74.51 -0.8109
-Llama2-GPT All 35.22 -0.7835 71.04 -0.7924
-Simple 33.9 -0.777 69.62 -0.7919
-Pro 35.45 -0.7916 71.54 -0.7881
-Mistral-Qwen All 44.9 -0.8722 80.82 -0.8794
-Simple 38.47 -0.8494 74.04 -0.8271
-Pro 44.32 -0.8045 80.08 -0.8682
-Mistral-GPT All 40.37 -0.8042 76.58 -0.8736
-Simple 35.51 -0.8297 72.27 -0.8664
-Pro 38.52 -0.7103 74.91 -0.7969
-Vicuna-Qwen All 42.94 -0.8771 79.23 -0.8365
-Simple 37.86 -0.758 73.85 -0.7168
-Pro 42.01 -0.8609 77.86 -0.886
-Vicuna-GPT All 38.69 -0.8621 74.83 -0.8672
-Simple 34.77 -0.8096 70.71 -0.7775
-Pro 37.37 -0.7794 73.98 -0.8728
-Table 3: The ratio and the Pearson-correlation between the $F_{c}$ and $F_{m}$ of the MCQs within the upper right and lower left $2\times 2$ and $3\times 3$ squares.
-For LLMs, ‘Llama2-Qwen’ refers that the $F_{c}$ and $F_{m}$ are calculated with Llama2 based on the Qwen-extracted triplet, and similarly hereinafter.
-For the Dataset column, ‘All’ stands for all the qualified MCQs after the triplet extraction, ‘Pro’ refers to the qualified MCQs that are the members of the mmlupro dataset while ‘Simple’ refers to the rest of the MCQs that are relatively easier.
-Appendix D Human Annotation
-As there is potential risk for the LLM-based Knowledge Preserving evaluation in TrinEval procedure (line 4 and line 10 in 1) that the API-based LLMs might still be able to answer the questions without sufficient knowledge since this is still prompting LLMs who have been trained on these datasets and “know” the original content, human annotation is also applied.
-The annotation of each MCQ encompasses three subtasks: (1) answering the question in the re-organized form, (2) answering the question in the original form, and (3) verdicting if the reformulation is Knowledge Preserving or not.
-For efficient annotation, we implemented a stratified sampling procedure by selecting one MCQ per subject from all 56 MMLU subjects (as a temporary compromise for limited time, which will be expanded later) under both Qwen and GPT reorganization paradigms.
-This yielded 112 representative questions (2 systems × 56 subjects) for evaluation.
-Three human annotators independently performed dual-form assessments through: (1) Direct question answering with the reformed format first and the original format; (2) Knowledge Preservation (K.P.) scoring across two dimensions: i.
-Knowledge adequacy (sufficiency for accurate response), ii.
-removal of redundant content using a 5-point scale (1=unsatisfactory; 2=major information are missed or unnecessary information is incorporated, but part is still acceptable, 3=need to take some time to understand, but can still solve the MCQ, 4=an element properly belonging to one triplet component appears in another, but does not impact MCQ solving; 5=optimal).
-We use a continuous rather than binary metric to mitigate the cognitive difference of the threshold between the annotators.
-Inter-rater reliability was ensured through consensus-building discussions prior to formal annotation.
-Final scores were aggregated using mean values to further mitigate individual annotator bias.
-The results are shown in Tab.
-4 below.
-Model TrinEval correct only both correct original correct only K.P. score
-Qwen 0.0% 96.296% 3.704% 4.101
-GPT 0.0% 96.667% 3.333% 4.369
-Table 4: Result of Human Annotation on LLM-based knowledge preserving (K.P.) evaluation in TrinEval
-Notably, our analysis reveals that over 95% of correctly answered MCQs maintained consistency across both original and paraphrased formats.
-Furthermore, human annotators rated our paraphrased questions mean K.P.
-scores exceeding 4.0 (on a 5-point scale), which means that the reformulated MCQs only somehow influence the readability of humans but do not impact the solvability of the original format.
-This provides empirical validation that our proposed TrinEval methodology effectively preserves necessary knowledge elements from original MCQ formulations, while the influence of the LLM memorization during the evaluation is rather limited.
-Experimental results under human annotation also reveal that Qwen underperforms GPT in key metrics, particularly in processing long-context texts where it occasionally omits background information (evidenced by excessive "N/A" assignments in the Context fields).
-This capability gap is further reflected in MCQ annotations: there are only MCQs that are merely correctly answered with the original format except for the correct MCQs with both formats.
-Appendix E Elaboration on dataset-related information
-We suppose that unintentional data contamination arises from crawling dataset pages (e.g., Hugging Face) during the compilation of LLM pretraining datasets.
-When researchers are organizing the pretraining corpus, one or more neighboring original data samples would be truncated and concatenated sequentially into a pretraining sample.
-Thus, according to Carlini’s theory Carlini et al.
-[2022a] and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
-Besides, the previous study Bordt et al.
-[2024] also applies a similar method, the “Header Completion Test”, for tabular data memorization detection.
-By offering the heading rows within a CSV file, they also find that providing the preceding data samples can help to detect the memorized dataset by LLMs, which also practically proves that offering samples within the same dataset would affect the memorization evocation.
-Following this, similar to the dataset name, we take the samples within the same dataset as the few-shot prompt in order to find out if the TrinEval re-organization method can avoid such memorization evocation phenomenon.
-Still, in Fig.
-4, we can see that the blue curves remain below the green ones for the “def-fsp” setting, which proves that TrinEval can restrain the memorization evocation.
-Appendix F Embedding distance of memorized and non-memorized MCQs
-As there are 57 different subjects within the MMLU dataset, we believe unrelated sequences would lead to increased embedding distance even though they are among the mastered knowledge points.
-Here, we try to filter out the unrelated samples for each sample and thus filter the closest samples at the $21e-$ 2 level in order to make sure there are not too many unrelated samples incorporated.
-To highlight this result more prominently, we employed a relatively stringent data filtering strategy in the paper and made it 1% of the closest samples in Section 4.4.
-In the following version, we will add this clarification part in the paper.
-Here in order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5%.
-The results are shown in Tab.
-5 (RM stands for rote memorization, and GCL stands for genuine capability learning).
-{tblr}
-cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, vline2-3,5 = 1, vline2,4,6 = 2-8, hline1-3,9 = -, Threshold & 1% 3% 5%
-Subset RM GCL RM GCL RM GCL
-Llama-Qwen 29.538 16.501 32.718 17.281 33.790 18.060
-Llama-GPT 28.925 17.777 31.327 18.275 32.315 19.068
-Mistral-Qwen 78.663 47.648 86.003 52.995 89.598 55.525
-Mistral-GPT 82.932 56.375 90.597 62.112 94.384 64.820
-Vicuna-Qwen 38.761 21.006 41.396 23.189 42.490 24.185
-Vicuna-GPT 37.037 22.996 41.396 23.189 42.490 24.185
-Table 5: Result of averaged embedding distance of the closest MCQs under different threshold
-We can see that, as we said above, the more samples we incorporated, the higher the average distance of the closest embeddings grows.
-Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings.
-This proves that the reported results are robust and solid.
+In this figure, we take each unique $F_{m}$ as the threshold to separate the qualified MCQs into the Memorized and non-Memorized MCQs.
+We compute the probability of a randomly selected non-Memorized MCQ’s $F_{c}$ exceeding a randomly selected Memorized MCQ’s $F_{c}$ under each threshold as the blue curve, and the green curve is the p-value of the T-test between the $F_{c}$s of the non-Memorized MCQs and the Memorized MCQs.